### v1

# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

Abstract
While recent multilingual automatic speech recognition models claim to support thousands of languages, ASR for low-resource languages remains highly unreliable due to limited bimodal speech and text training data. Better multilingual spoken language understanding (SLU) can strengthen massively the robustness of multilingual ASR by levering language semantics to compensate for scarce training data, such as disambiguating utterances via context or exploiting semantic similarities across languages. Even more so, SLU is indispensable for inclusive speech technology in roughly half of all living languages that lack a formal writing system. However, the evaluation of multilingual SLU remains limited to shallower tasks such as intent classification or language identification. To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses topical speech classification in 102 languages and multiple-choice question answering through listening comprehension in 92 languages. We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU. Our results show that cascaded systems exhibit greater robustness in multilingual SLU tasks, though speech encoders can achieve competitive performance in topical speech classification when appropriately pre-trained. We further find a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, highlighting the mutual benefits between acoustic and semantic speech representations.111The datasets are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
Fleurs-SLU: A Massively Multilingual Benchmark for
Spoken Language Understanding
Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila - Quebec AI Institute {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
1 Introduction
State-of-the-art multilingual automatic speech recognition (ASR) systems claim to support thousands of languages Pratap et al. (2024); Chen et al. (2024). However, these systems still struggle to reliably transcribe low-resource languages due to the scarcity of bimodal speech and text data in those languages.
Massively multilingual ASR models are commonly pre-trained on the self-supervised wav2vec-BERT 2.0 objective Baevski et al. (2020). Such pre-training learns acoustic features that align representations with phonetics rather than semantics of spoken language Choi et al. (2024).
We therefore argue that speech models require strong multilingual spoken language understanding (SLU) capabilities to enhance massively multilingual ASR for low-resource languages.
Multilingual speech models better at semantic encoding of speech tend to be more robust in ASR too, as they utilize context-related cues, leverage cross-lingual similarities (e.g., cognates, shared subwords, or similar syntax), and better support code-switching in mixed-language speech (cf. §5). Therefore, multilingual SLU capabilities are a key feature of speech models that enable sample-efficient cross-lingual transfer to low-resource or entirely unsupported languages.
Moreover, SLU is a crucial component of inclusive speech technology for languages without a formal writing system, as these languages cannot rely on coupling ASR with language models to offload semantic understanding. This limitation affects approximately half of all living languages.
Despite all this, existing multilingual datasets for evaluating the SLU of speech models have significant limitations. The Minds14 benchmark assesses SLU on 14 high-resource languages for intent classification, which often only requires identifying shallow semantics, such as detecting specific keywords Gerz et al. (2021). SpeechTaxi comprises spoken topical classification for Bible verses in 28 diverse languages, and as such does poorly corresponds to real-world usage scenarios Keller and Glavaš (2024).
To address the shortcomings of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores). We first filter out silent and noisy instances from Fleurs and then map the remaining data back to Flores Team et al. (2022).
The resulting dataset is then merged with SIB-200 Adelani et al. (2024) and Belebele Bandarkar et al. (2023) to create a topical utterance classification benchmark for 102 languages and a multiple-choice question answering on spoken paragraphs for 92 languages, respectively.
Contributions. 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mse). We extensively benchmark state-of-the-art speech models on multilingual semantic speech classification and multiple-choice from spoken paragraphs. Like prior work, we observe that Cascaded Systems (cs) remain the more robust option Keller and Glavaš (2024). We are the first to show that mse pre-trained on language understanding objectives can yield performance competitive to that of cs on topical speech classification. We further isolate utterance quality as a factor in SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can non-negligibly deteriorate transfer performance. 2) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT). This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust. 3) We ablate speech classification on SeamlessM4Tv2-Large Seamless Communication et al. (2023), comparing the performance of the model with and without the final length adaptor layer. We observe that performance decreases only slightly when accounting for the reduction in model size. This ablation further underscores that pre-training objectives, rather than architectural choices, drive multilingual SLU and, through SLU, the overall capabilities of multilingual speech technology.
2 Related Work
Multilingual Speech Representation Learning.
Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training. Among these models, mHubert employs the self-supervised wav2vec-BERT 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features Boito et al. (2024). MMS-1b is also pretrained with wav2vec-BERT 2.0 on an large corpus of 491K hours of speech spanning 1,406 languages Pratap et al. (2024). Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and VAD on 680k hours of audio Radford et al. (2022). SeamlessM4Tv2-Large is a multilingual multimodal translation model Seamless Communication et al. (2023). It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with wav2vec-BERT 2.0 on 4.5M hours of audio. The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives. The authors perform KD from the text encoder to the speech encoder by minimizing the KL-divergence between the decoder’s token output distributions on bi-modal speech and text data.
Multilingual SLU. The evaluation of multilingual Speech Language Understanding (SLU) is constrained to a limited set of tasks. SLU has been significantly shaped by task-oriented dialogue (ToD), leading to the creation of datasets focused on ToD-specific tasks such as intent classification and slot filling. These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords. Additionally, commonly used utterance-level SLU tasks like language identification and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encoding. As a result, the majority of SLU datasets are predominantly in English. Multilingual exceptions are otherwise limited. The Minds14 dataset for intent classification only includes 14 high-resource languages Gerz et al. (2021). SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains Keller and Glavaš (2024).
Concurrent Work. Costa-jussà et al. (2024) concurrently released the multilingual speech benchmark 2M-Belebele for 74 languages. In this dataset, the authors first extend Flores by ca. 20% by incorporating human recordings for both sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele. Costa-jussà et al. (2024) find that a cs-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by ca. 8 percentage points (pp).
3 Fleurs-SLU
In Fleurs-SLU, we postprocess and joins datasets that are based on Flores to create a massively multilingual SLU benchmark for speech classification and multiple-choice question answering from spoken paragraphs.
3.1 Core Datasets
Flores. Flores is a machine translation benchmark consisting of 3,001 sentences extracted from English Wikipedia paragraphs, which have been professionally translated into over 200 languages Team et al. (2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets. The authors did not release the test set.
Fleurs. Fleurs comprises 2.3 spoken utterances, on average, per sentence from the public dev and devtest splits of 102 languages in Flores Conneau et al. (2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores. Fleurs is used to evaluate multilingual ASR, language identification, as well as speech-to-text and text-to-speech translation in all language directions.
SIB-200. Adelani et al. (2024) refined the topical metadata annotations of sentences in the dev and devtest splits of Flores into 7 categories.444The categories are ”science/technology”, ”travel”, ”politics”, ”sports”, ”health”, ”entertainment”, ”geography”. The resulting SIB-200 is a natural language understanding benchmark for 205 language variants.
Belebele. Belebele is a multiple-choice machine reading comprehension benchmark parallel across 122 language variants Bandarkar et al. (2024). The authors first reconstruct paragraphs from Flores sentences. Four multiple-choice questions are created per English paragraph and then professionally translated into 121 language variants. Belebele comprises 900 questions, each associated with one of 488 distinct passages.
3.2 Benchmark Creation
We compile Fleurs-SLU by carefully aligning data from the above benchmarks as follows.
1) Merging Fleurs & Flores. We begin by removing silent instances from Fleurs.555The numbers of removed examples are listed in Appendix A.6. We first normalize the loudness to an RMS target level of . We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if voice activity is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect. Lastly, we manually verify our approach on 50 randomly sampled predicted silent instances. We find only one borderline misclassified example, which is very noisy but still comprehensible. We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
2a) SIB-Fleurs. For each language, we pool instances from the training, validation, and test splits of SIB-200 Adelani et al. (2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before. The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset. This segmentation ensures compatibility for speech models that may be trained on ASR using the training set of Fleurs prior to SIB-Fleurs evaluation. This also ensures that the speakers in the training set are different from those in the validation and test sets. Our final dataset comprises a median of 728 instances for training, 70 for validation, and 174 for testing across 102 languages. Figure 1 displays the available samples by split per language.999The full data is available in Appendix A.2.
2b) Belebele-Fleurs. We merge our Fleurs-Flores sentences with Belebele paragraphs by intersecting the URLs of the texts. We discard all paragraphs that are not complete in Fleurs-Flores. We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters). Belebele-Fleurs contains, on average, 709 evaluation instances (with a median of 772) across 92 language variants. Figure 2 illustrates the number of samples recovered from the original 900 paragraphs by language variant.
4 Experimental Setup
4.1 Tasks and Languages
SIB-Fleurs. We train mse on the utterances and cs on the transcriptions of the English training set. For both mse and cs, we feed the sequence-level representation pooled from token or speech frame embeddings into a classification head. Roberta-Large uses the [CLS] token as a sequence-level embedding. All other models average the token or speech frame output embeddings.
Belebele-Fleurs. For cs, we fine-tune and validate models on the English training and development sets compiled by Bandarkar et al. (2024), respectively. We jointly embed the paragraph, question, and choices with text encoders. We then average the token representations of each choice and project the resulting embedding via head to a logit . We minimize the cross-entropy of the concatenated choice logits to the label choice.101010We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input. We leave such evaluation to future work.
4.2 Cross-Lingual Transfer Setups
We experiment on two commonly used cross-lingual transfer (xlt) paradigms. Zero-shot cross-lingual transfer (zs-xlt) and Translate-Test (TTest) allow us to evaluate xlt without requiring additional annotation for any target language. In zs-xlt, we first train a multilingual model on source-language data and then directly run inference on target-language instances. In TTest, the model is first fine-tuned on labeled source-language data. At test time, the target-language examples are translated to the source language prior to inference, which enables xlt with monolingual LLMs.
Speech Classification. We evaluate state-of-the-art mse for speech classification that differ in pre-training objectives and training data (cf. §2). We include MMS-1b without intermediate fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on a broad range of multilingual datasets (‘MMS-1b-all’). For ASR fine-tuning, Pratap et al. (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen. For task fine-tuning, we freeze the language adapters to facilitate zs-xlt. We further evaluate on the speech encoders of Whisper-v3-Large and SeamlessM4Tv2-Large.
Cascading. Cascaded systems (cs) perform xlt in two steps. cs first transcribe speech into text with an ASR model. The transcription is then fed into an text encoder paired with a classification head. We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large. Note that speech-to-text-translation into English (S2ETT) with cs is equivalent to TTest as we merely monolingual LLMs are needed for xlt. We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024). For SIB-Fleurs, we base our cs experiments on Roberta-Large Liu et al. (2019).111111We did not find improvements by using LLMs in SIB-Fleurs and therefore run broad-spectrum experiments with the more efficient Roberta-Large. For Belebele-Fleurs, we run our experiments on LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024). LLM2Vec BehnamGhader et al. (2024) enables bidirectional attention and continues training on both masked next-token prediction and SimCSE Gao et al. (2021). NLLB-LLM2Vec fuses LLM2Vec with the encoder of the multilingual NLLB translation model to extend its NLU capabilities to more than 200 languages Schmidt et al. (2024).
4.3 Training Setup
Training Details. We train all models with the AdamW (Loshchilov and Hutter, 2019), weight decay of , and with 10% linear warm-up and then linear decay on an effective batch size of 32. For SIB-Fleurs, we run a grid search over the learning rates , since suitable hyperparameters have not yet been extensively studied for mse on such a downstream task. For Belebele-Fleurs, we fine-tune the LLMs on the learning rates with LoRAs of rank and alpha attached onto all linear layers, as full fine-tuning is prohibitively expensive. We train models for 20 and 3 epochs for SIB-Fleurs and Belebele-Fleurs, respectively, and validate at every 10% of training steps. Experimental results are averaged across 3 random seeds. We report results for runs on the learning rate that performs best, on average, on the English validation sets.
Utterance Quality. For each sentence, Fleurs comprises one or more utterances. We compute the character error rate (CER) between the ground truth sentence and its transcriptions.121212We compute CER against the original Flores sentence, as Flores represents the written form of the utterance best suited for processing by LLMs. We then fine-tune models on two sets of utterances: those with the lowest CER (best quality) and those with the highest CER (worst quality). We thereby isolate the downstream impact of utterance quality, ensure a balanced training set, and compare mse to cs more fairly, which benefit less from mixing transcriptions of varying quality.
5 Results
In-Language ZS-XLT Utterance Quality Best Worst Best Worst Best Worst Best Worst Best Worst SIB-Fleurs EN (1) Whisper (84) SeamlessM4T (89) Unsupported (9) Not-EN (101) MSE MMS-1b 44.6 46.9 19.9 13.6 19.8 13.5 18.1 11.3 19.8 13.4 MMS-1b-Fleurs 64.8 64.2 25.1 20.7 24.6 20.2 21.5 17.2 25.0 20.6 MMS-1B-All 55.2 53.1 21.6 18.1 21.4 17.9 19.9 16.8 21.7 18.2 Whisper-v3-L 78.3 76.3 45.9 41.9 45.3 41.2 35.9 31.1 44.7 40.8 SeamlessM4Tv2-L 87.4 88.5 82.3 79.3 82.2 79.1 65.6 61.0 79.0 75.8 CS Roberta-L-Wh-EN 90.4 91.7 75.7 72.6 73.5 70.6 45.8 42.7 71.5 68.4 Roberta-L-S4T-EN 91.7 90.8 86.1 84.5 85.9 84.2 65.9 63.7 82.3 80.6 Belebele-Fleurs EN (1) Whisper (78) SeamlessM4T (83) Unsupported (8) Not-EN (91) CS LLM2Vec-Wh-EN 94.7 94.4 60.1 58.9 58.5 57.4 32.1 32.4 56.5 55.5 LLM2Vec-S4T-EN 95.4 94.3 67.8 67.0 66.5 65.7 39.3 38.6 63.8 63.0 NLLB-LLM2Vec-Wh 94.9 93.4 57.9 56.0 56.6 54.9 37.4 36.6 55.4 53.8 NLLB-LLM2Vec-S4T 94.9 93.4 61.3 60.2 60.8 59.5 45.2 42.8 59.2 57.8
Table 1 summarizes the main results for both tasks by approach across our xlt setups. We dissect the results along several axes of analysis.
English. The first column presents the in-language (i.e., no-xlt) English performance by task and utterance quality for our models.
CS. English ASR performance is strong for both transcription models (cf. Figure 3), and cs effectively leverages the NLU capabilities of LLMs, thanks to their extensive pre-training on both tasks. As a result, cs outperforms all mse models on SIB-Fleurs. On Belebele-Fleurs, both cs models backed by LLM2Vec and NLLB-LLM2Vec perform comparably. In sum, the transcription model appears to have little to no impact on English performance.
MSE. The English results on SIB-Fleurs suggest that the SLU capabilities of mse models are strongly shaped by their pre-training curriculum. MMS-1b, which is pre-trained solely in a self-supervised manner, underperforms other mse on SIB-Fleurs (44.6%). Post-hoc fine-tuning of MMS-1b on ASR, either on Fleurs (MMS-1B-Fleurs) or Fleurs combined with additional data (MMS-1b-All), results in significant improvements on SIB-Fleurs of +10-20 percentage points (pp). Notably, MMS-1B-Fleurs outperforms MMS-1b-All (+9.6 pp), suggesting that the broader domain mixture in the MMS-1b-All training set negatively affects this downstream task. Whisper-v3-Large surpasses all MMS-1B variants (78.3%). The large-scale pre-training of Whisper-v3 on both multilingual ASR and S2ETT enhances its SLU performance. SeamlessM4Tv2-Large is the best performing mse on the English test set of SIB-Fleurs (87.4%), with only a slight performance gap compared to cs (ca. -3 pp). The joint pre-training on multilingual ASR as well as multilingual and cross-modal MT with text-to-speech knowledge distillation significantly enriches the semantics in speech representations of SeamlessM4Tv2-Large. The results particularly highlight that (i) multilingual ASR, (ii) S2ETT pre-training, and (iii) text-to-speech distillation are crucial training objectives for enabling speech encoders to acquire strong SLU capabilities.
Utterance Quality. The quality of English utterances does not seem to affect downstream performance for either mse or cs across both tasks. This robustness can be attributed to the exhaustive training of all backbone models on English ASR, supported by the large-scale availability of diverse English speech data.
zs-xlt. The right-hand side of Table 1 reports zs-xlt performance in various setups. For each task, we group the languages into (i) languages supported by Whisper-v3, (ii) languages supported by SeamlessM4Tv2, and (iii) languages unsupported by either model. The sets of languages of Fleurs that Whisper-v3 and SeamlessM4Tv2 support overlap in 81 languages.131313For unsupported languages, we hand-select the closest supported language to transcribe into.
CS. Across both tasks and all language groups, models trained on transcriptions of SeamlessM4Tv2-Large consistently outperform models fine-tuned on transcriptions of Whisper-v3-Large. The performance gap grows as Whisper’s support decreases or becomes unavailable for the target languages (cf. ‘SeamlessM4Tv2’ and ‘unsupported’ language groups in Table 1). For SIB-Fleurs, the cs of SeamlessM4Tv2-Large and Roberta-Large deteriorates only slightly in xlt performance to all 101 target language, on average, relative to English (ca. -9.6 pp). Most of this xlt gap comes from the 9 target languages that SeamlessM4Tv2 does not support (-25.6 pp on average). In contrast, pairing Roberta-Large with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (-10.4 pp) and a much larger deficit on languages neither model supports (-44.6 pp). For Belebele-Fleurs, S2ETT paired with LLM2Vec (i.e., TTest) outperforms zs-xlt cascading on in-language transcriptions and NLLB-LLM2Vec, except for unsupported languages. This suggests that S2ETT of both Whisper-v3 and SeamlesseM4Tv2 sufficiently translates the target languages into English for successful NLU. On unsupported languages, however, NLLB-LLM2Vec performs better for likely two reasons. First, for languages with low S2ETT quality, in-language transcriptions to closely related languages likely better preserve the core meanings of input sequences. Second, in addition to translation, NLLB was pre-trained with denoising autoencoding, making NLLB-LLM2Vec more resistant to noisy inputs than LLM2Vec. The findings suggests that SeamlessM4Tv2-Large is a much more robust transcription model for both in-language ASR as well as S2ETT (cf. Figure 3).
MSE. The zs-xlt performance of mse on SIB-Fleurs mirrors the trends we observed for English performance. When trained solely with the wav2vec-BERT 2.0 objective, MMS-1B (19%) only slightly surpasses random performance (14%) in zs-xlt for the highest-quality utterances, regardless of the target language group. In contrast to English, post-hoc modular ASR fine-tuning of MMS-1B with language adapters and language-specific decoding heads, whether on Fleurs alone (MMS-1B-Fleurs) or on Fleurs with additional data (MMS-1B-All), yields much smaller gains in zs-xlt (ca. +1-3 pp). Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English (i.e., in-language) performance to other languages effectively, with the xlt gap ranging from -32.4 pp for supported languages to -42.4 pp for unseen languages. In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp) while maintaining reasonable performance on unsupported languages (65.9%). These findings indicate that multilingual cross-modal machine translation and multilingual text-to-speech token distillation align and semantically enrich multilingual speech representations to significantly enhance cross-lingual SLU.
Utterance Quality. The utterance quality consistently affects performance on both tasks for all model configurations in zs-xlt. Notably, cs seem to be less affected than mse. We attribute this observation to the much more sizable pre-training on text of diverse quality of LLMs compared to speech models. For SIB-Fleurs, the magnitude of the performance gaps between best and worst quality utterances in zs-xlt mimics how well the speech model backbones perform across both tasks in both cs and end-to-end speech classification. SeamlessM4Tv2-Large is more robust to noisy utterances than Whisper-v3-Large, whereas MMS variants suffer from the largest drops (ca. 5 pp).
5.1 Further Analyses and Discussion
ASR & Translation Performance. To further understand the underlying factors behind our main results (cf. Table 1), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages. We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.141414We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large. We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model is highly comparable across splits. We trust that neither of the models has been trained on the dev and test set of Fleurs. We use the implementations for CER and sacreBELU of Huggingface’ evaluate. We then ‘macro-average’ the metrics over all languages. Figure 3 charts the results of this analysis.
ASR. Whisper-v3-Large outperforms SeamlessM4Tv2-Large for ASR on 58 out of 102 Fleurs languages (cf. ‘win ratio’). Nevertheless, SeamlessM4Tv2-Large is the overall more robust transcription model: in transcription for languages other than English, the CER of Whisper-v3-large is about twice as high as the CER of SeamlessM4Tv2-Large, on average. The full per-language results (cf. Appendix A.5)) confirm that SeamlessM4Tv2-Large exhibits much better ASR support for low(er)-resource languages, while being competitive in transcription quality for higher-resource languages.
S2ETT. SeamlessM4Tv2-Large outperforms Whisper-v3-large across the board in speech-to-English-text-translation and is favored for all but 2 languages. For both groups of languages supported by, Whisper-v3-Large and SeamlessM4Tv2-Large, respectively, SeamlessM4Tv2-Large achieves on average about 8 higher sacreBLEU than Whisper-v3-Large. The gap from SeamlessM4T-Large to Whisper-v3-Large reduces to about 5 sacreBLEU for unsupported languages. This supports the notion that SeamlessM4Tv2-Large is a more robust multilingual speech encoder.
The more robust ASR and much stronger translation performance of SeamlessM4Tv2-Large results stem from its pre-training. SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB-200 translation model Team et al. (2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on wav2vec-BERT 2.0 objective. The model is then trained on translation objectives from text and speech to text between any two languages in both translation directions. The text encoder-decoder backbone is used to train the speech encoder with token-level knowledge distillation objectives on decoder output representations. On the contrary, Whisper-v3 trains models from scratch on, among others, in-language ASR and S2ETT (cf. §4). Consequently, the mixture of strong initialization from existing MT backbones, text-to-speech knowledge distillation, and multimodal translation objectives result in much stronger translation performance.
The main results (cf. Table 1), together with our ASR and S2ETT analyses, underscore that sizable text and pre-training coupled with cross-modal and multilingual translation and text-to-speech knowledge distillation infuses rich semantic knowledge into mse, as witnessed by the SLU performance of SeamlessM4Tv2-based models on Fleurs-SLU.
Length Adaptor. The sequence length of encoded speech typically far exceeds the sequence length of embedded tokenized text for the same input. The mse of SeamlessM4Tv2-Large therefore appends a temporal convolution, called the length adaptor, as its final layer to reduce the resolution of speech frames by a factor of 8 Zhao et al. (2022); Seamless Communication et al. (2023).151515For architectural details on the length adaptor, we refer to the original paper Communication et al. (2023). For bimodal input, the length adaptor thereby mitigates the discrepancy in sequence lengths between speech encodings and tokenized text embeddings, intending to better align representations across modalities. This raises the question of whether a length adaptor is essential for pooling semantic meaning from speech output tokens. To investigate this, we ablate SeamlessM4Tv2-Large’s performance on SIB-Fleurs, comparing results with and without the length adaptor.161616We run a full grid search over learning rates (cf. §4) for SeamlessM4Tv2-Large without the length adaptor.
Table 2 presents the results for both variants. The results are mixed. For English, performance improves on high-quality utterances but decreases on lower-quality ones. In contrast, zs-xlt performance shows a slight but consistent decline, which is more pronounced for lower-quality utterances in supported languages and, irrespective of utterance quality, in unsupported languages. Three factors may explain this modest gap. First, removing the length adaptor reduces model capacity by 47M parameters (-7.4 pp), which were subjected to substantial pre-training. Second, the length adaptor, as the final layer of the speech encoder, is most explicitly trained to embed the input into a shared multilingual representation space, attended to by the text decoder at every layer. Finally, the length adaptor may improve the robustness of speech encodings by filtering noisy frames through temporal downsampling. While a modest gap persists across setups, we however cannot decisively infer based on the results whether appending a length adaptor is, as opposed to replacing model capacity with other layers of equal parameter size during pre-training, crucial for improved multilingual SLU. We conclude that the pre-training regime is more vital for multilingual SLU abilities of mse than nuanced architectural design choices.
ZS-XLT EN (1) S4T (89) N/A (12) AVG (101) S4Tv2-L Size B W B W B W B W Incl. LA 635M 87.4 88.5 82.2 79.1 55.0 51.4 79.0 75.8 Excl. LA 588M 88.5 86.4 81.0 77.6 52.7 49.1 77.8 74.4 47M +1.1 -2.1 -1.2 -2.5 -2.3 -2.3 -1.2 -1.4
6 Conclusion
We introduce Fleurs-SLU, a multilingual SLU benchmark for semantic speech classification across 102 languages and multiple-choice question answering from spoken paragraphs in 92 languages. Using Fleurs-SLU, we evaluate massively multilingual speech models in both end-to-end speech classification and a cascaded approach that combines initial speech-to-text transcription and subsequent text-based classification with LLMs. Our findings indicate that, while cascaded systems remain the most robust option, multilingual speech encoders can achieve competitive performance when adequately pre-trained. Furthermore, we observe a strong correlation between strong multilingual SLU and both the robustness of multilingual ASR and the effectiveness of cross-modal speech translation to English text. This suggests that multilingual SLU and multilingual ASR can be mutually beneficial. We hope that our findings inspire future work towards developing more efficient multilingual speech encoders that are jointly pre-trained for both multilingual ASR and SLU to close the performance gap between end-to-end speech classification and cascaded approaches.
Appendix A Appendix
A.1 Code
The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
A.2 SIB-Fleurs
Language Train Validation Test afr_Latn 406 86 95 amh_Ethi 752 54 149 arb_Arab 579 64 133 asm_Beng 730 71 176 ast_Latn 701 69 177 azj_Latn 712 71 174 bel_Cyrl 690 71 177 bul_Cyrl 749 70 176 ben_Beng 742 71 176 bos_Latn 746 71 177 cat_Latn 683 71 177 ceb_Latn 741 61 149 ckb_Arab 738 70 176 zho_Hans 751 71 176 ces_Latn 732 68 172 cym_Latn 739 71 177 dan_Latn 696 70 177 deu_Latn 736 69 175 ell_Grek 750 67 168 eng_Latn 738 71 177 spa_Latn 676 71 177 est_Latn 700 71 176 pes_Arab 692 66 165 fin_Latn 735 71 175 tgl_Latn 604 71 176 fra_Latn 753 65 164 gle_Latn 731 71 176 glg_Latn 660 71 174 guj_Gujr 752 71 177 hau_Latn 753 70 166 heb_Hebr 754 70 175 hin_Deva 653 60 132 hrv_Latn 756 71 176 hun_Latn 750 71 177 hye_Armn 741 71 177 ind_Latn 728 69 167 ibo_Latn 737 71 177 isl_Latn 381 18 23 ita_Latn 743 69 175 jpn_Jpan 662 62 164 jav_Latn 740 67 171 kat_Geor 557 69 177 kam_Latn 752 69 179 kea_Latn 725 71 175 kaz_Cyrl 749 70 176 khm_Khmr 588 69 168 kan_Knda 660 70 174 kor_Hang 669 61 141 kir_Cyrl 729 71 177 ltz_Latn 703 71 176 lug_Latn 691 70 173
Language Train Validation Test lin_Latn 755 59 139 lao_Laoo 591 54 132 lit_Latn 730 71 178 luo_Latn 698 39 98 lvs_Latn 634 69 174 mri_Latn 749 71 176 mkd_Cyrl 680 71 177 mal_Mlym 723 68 174 khk_Cyrl 743 71 177 mar_Deva 749 71 177 zsm_Latn 713 67 171 mlt_Latn 731 71 176 mya_Mymr 746 71 175 nob_Latn 723 51 127 npi_Deva 754 70 175 nld_Latn 729 58 123 nso_Latn 633 70 169 nya_Latn 720 68 169 oci_Latn 756 71 177 ory_Orya 442 71 168 pan_Guru 580 56 143 pol_Latn 723 68 165 pbt_Arab 701 55 144 por_Latn 728 70 177 ron_Latn 734 69 177 rus_Cyrl 733 71 173 snd_Arab 749 71 177 slk_Latn 628 71 169 slv_Latn 704 71 174 sna_Latn 689 71 176 som_Latn 746 70 177 srp_Cyrl 730 63 164 swe_Latn 686 71 168 swh_Latn 745 65 154 tam_Taml 693 71 169 tel_Telu 658 66 153 tgk_Cyrl 680 69 163 tha_Thai 710 71 176 tur_Latn 692 67 164 ukr_Cyrl 732 67 164 umb_Latn 473 39 108 urd_Arab 636 65 120 uzn_Latn 734 69 175 vie_Latn 737 70 176 wol_Latn 643 52 123 xho_Latn 756 71 177 yor_Latn 686 71 172 zho_Hant 624 70 172 zul_Latn 739 69 175 fuv_Latn 752 68 166 gaz_Latn 574 6 17
A.3 Belebele-Fleurs
Language Samples afr_Latn 309 amh_Ethi 782 arb_Arab 387 asm_Beng 824 azj_Latn 759 ben_Beng 855 bul_Cyrl 873 cat_Latn 652 ceb_Latn 783 ces_Latn 802 ckb_Arab 842 dan_Latn 696 deu_Latn 804 ell_Grek 837 eng_Latn 844 est_Latn 736 fin_Latn 826 fra_Latn 839 fuv_Latn 848 gaz_Latn 25 guj_Gujr 880 hau_Latn 838 heb_Hebr 878 hin_Deva 515 hrv_Latn 896 hun_Latn 879 hye_Armn 861 ibo_Latn 838 ind_Latn 783 isl_Latn 81 ita_Latn 851 jav_Latn 835 jpn_Jpan 590 kan_Knda 606 kat_Geor 372 kaz_Cyrl 870 kea_Latn 770 khk_Cyrl 869 khm_Khmr 439 kir_Cyrl 811 kor_Hang 535 lao_Laoo 346 lin_Latn 778 lit_Latn 834 lug_Latn 703 luo_Latn 512
Language Samples lvs_Latn 555 mal_Mlym 809 mar_Deva 869 mkd_Cyrl 667 mlt_Latn 816 mri_Latn 877 mya_Mymr 864 nld_Latn 674 nob_Latn 635 npi_Deva 876 nso_Latn 569 nya_Latn 752 ory_Orya 220 pan_Guru 396 pbt_Arab 628 pes_Arab 673 pol_Latn 765 por_Latn 791 ron_Latn 815 rus_Cyrl 819 slk_Latn 513 slv_Latn 724 sna_Latn 735 snd_Arab 878 som_Latn 874 spa_Latn 659 srp_Cyrl 766 swe_Latn 681 swh_Latn 780 tam_Taml 714 tel_Telu 567 tgk_Cyrl 632 tgl_Latn 505 tha_Thai 745 tur_Latn 706 ukr_Cyrl 773 urd_Arab 482 uzn_Latn 812 vie_Latn 847 wol_Latn 495 xho_Latn 900 yor_Latn 652 zho_Hans 888 zho_Hant 527 zsm_Latn 749 zul_Latn 838
A.4 Analysis of Speech-To-English-Text-Translation Performance
Language Whisper-v3-Large SeamlessM4Tv2-Large AVG 13.4 23.6 afr_Latn 32.9 42.5 amh_Ethi 0.8 18.5 arb_Arab 19.3 33.4 asm_Beng 2.3 21.9 ast_Latn 25.7 27.7 azj_Latn 10.7 18.3 bel_Cyrl 10.6 17.9 ben_Beng 8.1 27.2 bos_Latn 27.6 35.6 bul_Cyrl 26.4 33.2 cat_Latn 32.5 39.9 ceb_Latn 6.2 9.1 ces_Latn 24.5 33.1 ckb_Arab 1.6 23.2 cym_Latn 9.2 33.2 dan_Latn 32.6 38.9 deu_Latn 32.6 36.8 ell_Grek 20.5 27.7 est_Latn 15.3 30.4 fin_Latn 19.2 27.4 fra_Latn 34.0 36.1 fuv_Latn 0.2 0.9 gaz_Latn 0.3 0.7 gle_Latn 1.5 18.6 glg_Latn 27.9 35.6 guj_Gujr 10.7 30.2 hau_Latn 0.5 1.3 heb_Hebr 16.1 31.3 hin_Deva 19.3 28.5 hrv_Latn 24.9 31.7 hun_Latn 17.5 27.9 hye_Armn 9.0 31.2 ibo_Latn 0.6 2.6 ind_Latn 26.3 30.7 isl_Latn 7.5 25.7 ita_Latn 23.8 27.6 jav_Latn 4.1 23.6 jpn_Jpan 16.0 17.8 kam_Latn 0.8 2.5 kan_Knda 6.7 24.9 kat_Geor 2.3 21.7 kaz_Cyrl 3.8 25.0 kea_Latn 26.5 28.7 khk_Cyrl 0.8 18.5 khm_Khmr 4.5 22.3 kir_Cyrl 2.4 19.0 kor_Hang 19.0 22.7 lao_Laoo 7.1 25.9 lin_Latn 0.5 1.5 lit_Latn 12.7 24.3
Language Whisper-v3-Large SeamlessM4Tv2-Large ltz_Latn 16.0 17.6 lug_Latn 0.7 18.3 luo_Latn 0.9 1.3 lvs_Latn 13.3 29.0 mal_Mlym 9.9 25.0 mar_Deva 9.9 27.2 mkd_Cyrl 26.0 35.6 mlt_Latn 11.1 40.1 mri_Latn 6.5 1.3 mya_Mymr 0.4 19.4 nld_Latn 22.1 25.8 nob_Latn 29.7 34.7 npi_Deva 10.6 9.4 nso_Latn 0.7 2.6 nya_Latn 0.9 19.3 oci_Latn 17.6 23.5 ory_Orya 4.8 26.8 pan_Guru 12.9 30.2 pbt_Arab 1.7 18.2 pes_Arab 16.0 30.5 pol_Latn 20.7 24.2 por_Latn 37.5 38.9 ron_Latn 30.6 36.5 rus_Cyrl 26.2 30.3 slk_Latn 25.1 33.6 slv_Latn 18.0 27.4 sna_Latn 1.0 3.8 snd_Arab 3.6 8.7 som_Latn 0.4 18.1 spa_Latn 22.0 24.6 srp_Cyrl 29.8 37.3 swe_Latn 34.3 38.5 swh_Latn 5.6 32.4 tam_Taml 6.0 22.5 tel_Telu 10.3 26.5 tgk_Cyrl 8.7 26.9 tgl_Latn 21.2 25.6 tha_Thai 13.4 23.5 tur_Latn 22.2 30.1 ukr_Cyrl 27.6 32.8 umb_Latn 0.2 1.1 urd_Arab 15.9 25.5 uzn_Latn 5.0 25.5 vie_Latn 19.2 26.4 wol_Latn 1.2 1.6 xho_Latn 0.8 7.6 yor_Latn 0.6 14.6 zho_Hans 14.6 22.4 zho_Hant 8.7 18.5 zsm_Latn 24.7 31.0 zul_Latn 0.6 12.4
A.5 Analysis of ASR Performance
Language Whisper-v3-Large SeamlessM4Tv2-Large afr_Latn 13.1 12.6 amh_Ethi 207.8 23.9 arb_Arab 7.6 8.0 asm_Beng 99.0 15.1 ast_Latn 15.5 19.8 azj_Latn 7.2 8.3 bel_Cyrl 11.9 6.9 ben_Beng 33.9 9.0 bos_Latn 4.6 7.8 bul_Cyrl 5.2 9.6 cat_Latn 3.3 5.3 ceb_Latn 17.4 19.0 ces_Latn 4.2 8.0 ckb_Arab 98.3 13.9 cym_Latn 15.8 14.4 dan_Latn 4.8 10.0 deu_Latn 2.2 6.5 ell_Grek 6.4 10.8 eng_Latn 3.1 6.2 est_Latn 7.3 11.0 fin_Latn 3.2 10.2 fra_Latn 2.8 6.6 fuv_Latn 194.3 57.0 gaz_Latn 35.8 47.8 gle_Latn 85.1 21.4 glg_Latn 4.3 6.5 guj_Gujr 21.3 10.3 hau_Latn 33.8 51.8 heb_Hebr 12.2 14.5 hin_Deva 11.7 10.2 hrv_Latn 4.5 8.5 hun_Latn 5.2 8.5 hye_Armn 18.0 9.1 ibo_Latn 42.9 59.3 ind_Latn 4.0 9.2 isl_Latn 16.0 10.3 ita_Latn 2.1 5.2 jav_Latn 26.3 11.3 jpn_Jpan 6.3 16.4 kam_Latn 45.1 53.9 kan_Knda 20.1 10.9 kat_Geor 19.8 7.1 kaz_Cyrl 8.9 9.7 kea_Latn 35.8 38.1 khk_Cyrl 37.6 14.6 khm_Khmr 144.0 29.8 kir_Cyrl 28.3 8.8 kor_Hang 7.4 10.3 lao_Laoo 109.7 30.3 lin_Latn 22.0 58.7 lit_Latn 8.3 11.9 ltz_Latn 29.5 40.1
Language Whisper-v3-Large SeamlessM4Tv2-Large lug_Latn 44.9 14.1 luo_Latn 42.7 58.8 lvs_Latn 7.3 8.2 mal_Mlym 104.3 14.2 mar_Deva 24.3 10.7 mkd_Cyrl 5.8 8.6 mlt_Latn 26.0 12.1 mri_Latn 13.4 54.7 mya_Mymr 132.9 22.2 nld_Latn 3.1 6.8 nob_Latn 4.6 9.0 npi_Deva 25.6 74.1 nso_Latn 95.7 66.9 nya_Latn 35.6 12.4 oci_Latn 25.9 33.8 ory_Orya 93.8 11.4 pan_Guru 44.2 10.3 pbt_Arab 36.8 21.6 pes_Arab 8.6 7.7 pol_Latn 2.3 7.9 por_Latn 2.8 7.6 ron_Latn 3.2 6.6 rus_Cyrl 2.6 6.1 slk_Latn 3.9 7.0 slv_Latn 6.0 9.2 sna_Latn 26.7 32.7 snd_Arab 94.5 33.5 som_Latn 34.9 17.9 spa_Latn 2.2 6.3 srp_Cyrl 76.6 7.6 swe_Latn 3.7 9.2 swh_Latn 11.0 8.9 tam_Taml 13.3 11.6 tel_Telu 90.1 13.3 tgk_Cyrl 28.8 8.8 tgl_Latn 5.0 12.0 tha_Thai 9.8 11.4 tur_Latn 6.2 7.5 ukr_Cyrl 3.0 8.9 umb_Latn 156.7 55.5 urd_Arab 9.4 9.6 uzn_Latn 25.0 8.4 vie_Latn 4.7 6.7 wol_Latn 140.6 47.9 xho_Latn 48.1 32.3 yor_Latn 49.3 33.2 zho_Hans 16.2 19.0 zho_Hant 30.6 35.1 zsm_Latn 3.4 8.5 zul_Latn 42.6 18.4
A.6 Silent Fleurs Examples
Language Split Count nb_no train 497 es_419 train 490 cy_gb train 394 sd_in train 307 ny_mw train 15 ckb_iq train 8 ny_mw test 8 wo_sn train 7 nso_za test 6 ny_mw dev 6 ur_pk test 6 ps_af train 4 fa_ir train 4 so_so train 4 ceb_ph train 3 lg_ug train 3 kea_cv train 2 bg_bg train 2 bn_in train 2 cy_gb test 2 ff_sn train 2 hr_hr train 2 hy_am train 2 nso_za dev 2 ur_pk dev 2 ar_eg train 1 da_dk test 1 da_dk train 1 en_us train 1 ff_sn dev 1 ha_ng train 1 he_il dev 1 he_il test 1 he_il train 1 ig_ng train 1 kam_ke train 1 kn_in dev 1 kn_in test 1 kn_in train 1 mi_nz dev 1 mn_mn train 1 ms_my train 1 or_in train 1 sk_sk dev 1 sk_sk test 1 so_so test 1 ta_in train 1 te_in test 1 te_in train 1 umb_ao train 1

### v2

@@
 Abstract
-While recent multilingual automatic speech recognition models claim to support thousands of languages, ASR for low-resource languages remains highly unreliable due to limited bimodal speech and text training data.
-Better multilingual spoken language understanding (SLU) can strengthen massively the robustness of multilingual ASR by levering language semantics to compensate for scarce training data, such as disambiguating utterances via context or exploiting semantic similarities across languages.
-Even more so, SLU is indispensable for inclusive speech technology in roughly half of all living languages that lack a formal writing system.
-However, the evaluation of multilingual SLU remains limited to shallower tasks such as intent classification or language identification.
-To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses topical speech classification in 102 languages and multiple-choice question answering through listening comprehension in 92 languages.
+Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system, since these languages cannot pair automatic speech recognition (ASR) with language models to benefit from language technology.
+Even if low-resource languages possess a writing system, ASR for these languages remains unreliable due to limited bimodal speech and text training data.
+Better SLU can strengthen the robustness of massively multilingual ASR by levering language semantics to disambiguate utterances via context or exploiting semantic similarities across languages.
+However, the evaluation of multilingual SLU remains limited to shallow tasks such as intent classification or language identification.
+To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering through listening comprehension spanning 944 hours of speech across 92 languages.
 We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU.
@@
 Spoken Language Understanding
-Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila - Quebec AI Institute {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
+Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila, McGill University and Canada CIFAR AI Chair {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
 1 Introduction
-State-of-the-art multilingual automatic speech recognition (ASR) systems claim to support thousands of languages Pratap et al.
-(2024); Chen et al.
+Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology Ethnologue (2017).
+Spoken language understanding (SLU) is hence a critical feature of inclusive technology for these languages, which cannot rely on combining automatic speech recognition (ASR) with language models to offload semantic speech understanding.
+However, the datasets currently available for evaluating SLU in truly relevant languages suffer from significant limitations.
+For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords Gerz et al.
 (2021).
-However, these systems still struggle to reliably transcribe low-resource languages due to the scarcity of bimodal speech and text data in those languages.
-Massively multilingual ASR models are commonly pre-trained on the self-supervised wav2vec-BERT 2.0 objective Baevski et al.
+Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage Keller and Glavaš (2024).
+Moreover, even when a low-resource language has a writing system, multilingual ASR often struggles to reliably transcribe that language due to the limited availability of bimodal speech and text data.
+Pre-trained on the self-supervised wav2vec-BERT objective Baevski et al.
+(2020), these models embed acoustic features that capture phonetic rather than semantic information Choi et al.
 (2024).
-Such pre-training learns acoustic features that align representations with phonetics rather than semantics of spoken language Choi et al.
-(2024).
-We therefore argue that speech models require strong multilingual spoken language understanding (SLU) capabilities to enhance massively multilingual ASR for low-resource languages.
-Multilingual speech models better at semantic encoding of speech tend to be more robust in ASR too, as they utilize context-related cues, leverage cross-lingual similarities (e.g., cognates, shared subwords, or similar syntax), and better support code-switching in mixed-language speech (cf.
-§5).
-Therefore, multilingual SLU capabilities are a key feature of speech models that enable sample-efficient cross-lingual transfer to low-resource or entirely unsupported languages.
-Moreover, SLU is a crucial component of inclusive speech technology for languages without a formal writing system, as these languages cannot rely on coupling ASR with language models to offload semantic understanding.
-This limitation affects approximately half of all living languages.
-Despite all this, existing multilingual datasets for evaluating the SLU of speech models have significant limitations.
-The Minds14 benchmark assesses SLU on 14 high-resource languages for intent classification, which often only requires identifying shallow semantics, such as detecting specific keywords Gerz et al.
-(2021).
-SpeechTaxi comprises spoken topical classification for Bible verses in 28 diverse languages, and as such does poorly corresponds to real-world usage scenarios Keller and Glavaš (2024).
-To address the shortcomings of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores).
-We first filter out silent and noisy instances from Fleurs and then map the remaining data back to Flores Team et al.
+Conversely, multilingual speech models that excel in semantic encoding tend to be more robust in ASR tasks by, e.g., leveraging context-related cues or cross-lingual similarities such as cognates or similar syntactic structures (cf.
+§5.1).
+This underscores that robust SLU is an instrumental component of genuinely inclusive speech technology.
+To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores).
+We first filter out silent instances from Fleurs Conneau et al.
+(2022) and then map the remaining data back to Flores Team et al.
 (2022).
@@
 (2024) and Belebele Bandarkar et al.
-(2023) to create a topical utterance classification benchmark for 102 languages and a multiple-choice question answering on spoken paragraphs for 92 languages, respectively.
+(2023) to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
 Contributions.
 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mSE).
-We extensively benchmark state-of-the-art speech models on multilingual semantic speech classification and multiple-choice from spoken paragraphs.
-Like prior work, we observe that Cascaded Systems (cs) remain the more robust option Keller and Glavaš (2024).
-We are the first to show that mse pre-trained on language understanding objectives can yield performance competitive to that of cs on topical speech classification.
-We further isolate utterance quality as a factor in SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can non-negligibly deteriorate transfer performance.
-2) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT).
+SIB-Fleurs supports semantic utterance classification with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks.
+Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages Costa-jussà et al.
+(2024).
+2) We extensively benchmark state-of-the-art speech models on SIB-Fleurs and Belebele-Fleurs.
+Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding Keller and Glavaš (2024).
+We are the first to show that mSE pre-trained on language understanding objectives can yield performance competitive to that of CS on speech classification.
+We further isolate utterance quality as an important factor in multilingual SLU and show that, in zero-shot cross-lingual transfer (zs-xlt), worse utterance quality can substantially deteriorate transfer performance.
+3) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT).
 This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust.
-3) We ablate speech classification on SeamlessM4Tv2-Large Seamless Communication et al.
-(2023), comparing the performance of the model with and without the final length adaptor layer.
-We observe that performance decreases only slightly when accounting for the reduction in model size.
-This ablation further underscores that pre-training objectives, rather than architectural choices, drive multilingual SLU and, through SLU, the overall capabilities of multilingual speech technology.
 2 Related Work
@@
 (2024).
-Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and VAD on 680k hours of audio Radford et al.
+Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio Radford et al.
 (2022).
@@
 Multilingual SLU.
-The evaluation of multilingual Speech Language Understanding (SLU) is constrained to a limited set of tasks.
-SLU has been significantly shaped by task-oriented dialogue (ToD), leading to the creation of datasets focused on ToD-specific tasks such as intent classification and slot filling.
+The evaluation of multilingual SLU is constrained to a limited set of tasks.
+SLU has been significantly shaped by task-oriented dialogue (ToD), with datasets created focusing on ToD-specific tasks such as intent classification and slot filling.
 These tasks frequently require only basic semantic understanding, often reducing to merely detecting specific keywords.
-Additionally, commonly used utterance-level SLU tasks like language identification and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encoding.
+Additionally, commonly used utterance-level SLU tasks like language identification (LID) and sentiment classification do not assess content-based understanding but instead rely on phonetic or prosodic features of speech encodings.
 As a result, the majority of SLU datasets are predominantly in English.
@@
 Costa-jussà et al.
-(2024) concurrently released the multilingual speech benchmark 2M-Belebele for 74 languages.
+(2024) concurrently released the multilingual SLU benchmark 2M-Belebele for 74 languages.
 In this dataset, the authors first extend Flores by ca.
@@
 3 Fleurs-SLU
-In Fleurs-SLU, we postprocess and joins datasets that are based on Flores to create a massively multilingual SLU benchmark for speech classification and multiple-choice question answering from spoken paragraphs.
+We postprocess datasets that are based on Flores to create ‘Fleurs-SLU’, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs.
 3.1 Core Datasets
 Flores.
-Flores is a machine translation benchmark consisting of 3,001 sentences extracted from English Wikipedia paragraphs, which have been professionally translated into over 200 languages Team et al.
+The authors professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation Team et al.
 (2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets.
@@
 Fleurs.
-Fleurs comprises 2.3 spoken utterances, on average, per sentence from the public dev and devtest splits of 102 languages in Flores Conneau et al.
+Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores Conneau et al.
 (2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores.
-Fleurs is used to evaluate multilingual ASR, language identification, as well as speech-to-text and text-to-speech translation in all language directions.
+Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech translation in all language directions.
 SIB-200.
@@
 Belebele.
-Belebele is a multiple-choice machine reading comprehension benchmark parallel across 122 language variants Bandarkar et al.
+Belebele is a multiple-choice reading comprehension benchmark for 122 languages Bandarkar et al.
 (2024).
-The authors first reconstruct paragraphs from Flores sentences.
-Four multiple-choice questions are created per English paragraph and then professionally translated into 121 language variants.
-Belebele comprises 900 questions, each associated with one of 488 distinct passages.
+The authors reconstruct paragraphs from Flores sentences.
+They generate 1–2 questions per English paragraph, which are professionally translated into 121 languages.
+Belebele comprises 900 questions spanning 488 passages.
 3.2 Benchmark Creation
@@
 1) Merging Fleurs & Flores.
-We begin by removing silent instances from Fleurs.555The numbers of removed examples are listed in Appendix A.6.
+We begin by removing silent instances from Fleurs.555Counts of removed examples are listed in Appendix A.6.
 We first normalize the loudness to an RMS target level of .
-We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if voice activity is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect.
-Lastly, we manually verify our approach on 50 randomly sampled predicted silent instances.
+We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect.
+Lastly, we verify our approach on 50 randomly sampled predicted silent instances.
 We find only one borderline misclassified example, which is very noisy but still comprehensible.
@@
 This also ensures that the speakers in the training set are different from those in the validation and test sets.
-Our final dataset comprises a median of 728 instances for training, 70 for validation, and 174 for testing across 102 languages.
-Figure 1 displays the available samples by split per language.999The full data is available in Appendix A.2.
+Table 2 lists aggregated statistics on the instance- and the utterance-level for SIB-Fleurs.
+Appendix A.2 provides a list of samples by split per language.
+| Task | Utterance |
+| Classification | |
+| Languages | 102 |
+| Classes (Topics) | 7 |
+| Utterances per sample | 2.2 (2.0) |
+| Duration per utterance (s) | 12.9 (11.6) |
+| Total audio (hr) | 692 |
+| Samples by Split | |
+| Training | 696 (728) |
+| Validation | 66 (70) |
+| Test | 163 (174) |
 2b) Belebele-Fleurs.
@@
 We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters).
-Belebele-Fleurs contains, on average, 709 evaluation instances (with a median of 772) across 92 language variants.
-Figure 2 illustrates the number of samples recovered from the original 900 paragraphs by language variant.
+Table 2 provides a detailed summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
+Task Multiple-Choice QA Spoken Paragraphs Languages 92 Answer Choices 4 Questions per paragraph 1.8 (1.8) Sentences per paragraph 3.6 (4.0) Utterances per sentence 2.0 (2.0) Duration per utterance (s) 12.6 (12.0) Duration per paragraph (s) 47.0 (43.4) Total audio (hr) 944 Samples (Paragraphs) 709 (771)
 4 Experimental Setup
@@
 Belebele-Fleurs.
-For cs, we fine-tune and validate models on the English training and development sets compiled by Bandarkar et al.
+We train and validate models on the English training and dev sets of Bandarkar et al.
 (2024), respectively.
 We jointly embed the paragraph, question, and choices with text encoders.
-We then average the token representations of each choice and project the resulting embedding via head to a logit .
-We minimize the cross-entropy of the concatenated choice logits to the label choice.101010We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input.
+We then average the token encodings of each choice and project the choice embedding via head to a logit .
+We minimize the cross-entropy of the concatenated choice logits to the label choice.999We do not evaluate speech LLMs as existing models are based on Whisper-v3 which is limited to 30 seconds of audio input.
 We leave such evaluation to future work.
@@
 Speech Classification.
-We evaluate state-of-the-art mse for speech classification that differ in pre-training objectives and training data (cf.
+We evaluate state-of-the-art mSE for speech classification (cf.
 §2).
-We include MMS-1b without intermediate fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on a broad range of multilingual datasets (‘MMS-1b-all’).
+We include MMS-1b without fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1b-all’).
 For ASR fine-tuning, Pratap et al.
 (2024) train language adapters and language-specific decoding heads while keeping all other parameters frozen.
-For task fine-tuning, we freeze the language adapters to facilitate zs-xlt.
-We further evaluate on the speech encoders of Whisper-v3-Large and SeamlessM4Tv2-Large.
-Cascading.
+For task fine-tuning, we freeze the adapters to facilitate zs-xlt.
+We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large.
+Cascading & Text.
 Cascaded systems (CS) perform xlt in two steps.
@@
 We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large.
-Note that speech-to-text-translation into English (S2ETT) with cs is equivalent to TTest as we merely monolingual LLMs are needed for xlt.
+Note that speech-to-text-translation into English (S2ETT) with CS is equivalent to TTest as merely monolingual LLMs are needed for xlt.
+We further evaluate on the original text of SIB and Bebelebe, which we translate to English with SeamlessM4Tv2-Large.101010We use the gold Q&A for the respective language (i.e., English for ).
+Otherwise, we would need to interleave MT models, e.g., S2ETT of paragraphs with Whisper, but translating textual Q&A with SeamlessM4Tv2.
 We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024).
-For SIB-Fleurs, we base our cs experiments on Roberta-Large Liu et al.
-(2019).111111We did not find improvements by using LLMs in SIB-Fleurs and therefore run broad-spectrum experiments with the more efficient Roberta-Large.
-For Belebele-Fleurs, we run our experiments on LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024).
+We run our text-based and CS experiments on Roberta-Large Liu et al.
+(2019) as well as LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024).
 LLM2Vec BehnamGhader et al.
@@
 (2021).
-NLLB-LLM2Vec fuses LLM2Vec with the encoder of the multilingual NLLB translation model to extend its NLU capabilities to more than 200 languages Schmidt et al.
+NLLB-LLM2Vec fuses LLM2Vec with the encoder of the NLLB translation model to extend its NLU capabilities to 200 languages Schmidt et al.
 (2024).
@@
 Training Details.
-We train all models with the AdamW (Loshchilov and Hutter, 2019), weight decay of , and with 10% linear warm-up and then linear decay on an effective batch size of 32.
+We train all models with AdamW (Loshchilov and Hutter, 2019), weight decay of , and with 10% linear warm-up and followed by linear decay on an effective batch size of 32.
 For SIB-Fleurs, we run a grid search over the learning rates , since suitable hyperparameters have not yet been extensively studied for mSE on such a downstream task.
@@
 We report results for runs on the learning rate that performs best, on average, on the English validation sets.
+Text vs.
 Utterance Quality.
+We evaluate models trained on the original text (Text) as well as on varying quality of speech.
 For each sentence, Fleurs comprises one or more utterances.
-We compute the character error rate (CER) between the ground truth sentence and its transcriptions.121212We compute CER against the original Flores sentence, as Flores represents the written form of the utterance best suited for processing by LLMs.
+We compute the character error rate (CER) between the ground truth sentence and its transcriptions.111111We compute CER against the original Flores sentence, as Flores represents the written form of the utterance best suited for processing by LLMs.
 We then fine-tune models on two sets of utterances: those with the lowest CER (best quality) and those with the highest CER (worst quality).
@@
 5 Results
-In-Language ZS-XLT Utterance Quality Best Worst Best Worst Best Worst Best Worst Best Worst SIB-Fleurs EN (1) Whisper (84) SeamlessM4T (89) Unsupported (9) Not-EN (101) MSE MMS-1b 44.6 46.9 19.9 13.6 19.8 13.5 18.1 11.3 19.8 13.4 MMS-1b-Fleurs 64.8 64.2 25.1 20.7 24.6 20.2 21.5 17.2 25.0 20.6 MMS-1B-All 55.2 53.1 21.6 18.1 21.4 17.9 19.9 16.8 21.7 18.2 Whisper-v3-L 78.3 76.3 45.9 41.9 45.3 41.2 35.9 31.1 44.7 40.8 SeamlessM4Tv2-L 87.4 88.5 82.3 79.3 82.2 79.1 65.6 61.0 79.0 75.8 CS Roberta-L-Wh-EN 90.4 91.7 75.7 72.6 73.5 70.6 45.8 42.7 71.5 68.4 Roberta-L-S4T-EN 91.7 90.8 86.1 84.5 85.9 84.2 65.9 63.7 82.3 80.6 Belebele-Fleurs EN (1) Whisper (78) SeamlessM4T (83) Unsupported (8) Not-EN (91) CS LLM2Vec-Wh-EN 94.7 94.4 60.1 58.9 58.5 57.4 32.1 32.4 56.5 55.5 LLM2Vec-S4T-EN 95.4 94.3 67.8 67.0 66.5 65.7 39.3 38.6 63.8 63.0 NLLB-LLM2Vec-Wh 94.9 93.4 57.9 56.0 56.6 54.9 37.4 36.6 55.4 53.8 NLLB-LLM2Vec-S4T 94.9 93.4 61.3 60.2 60.8 59.5 45.2 42.8 59.2 57.8
+In-Language ZS-XLT Utterance Quality Best Worst Best Worst Best Worst Best Worst Best Worst SIB-Fleurs EN (1) Whisper (85) SeamlessM4T (90) Unsupported (7) Not-EN (101) mSE mHuBERT 41.1 39.7 27.3 19.1 27.2 19.1 26.5 16.3 26.9 18.6 MMS-1b 44.6 46.9 19.9 13.6 19.8 13.5 20.8 13.3 19.8 13.4 MMS-1b-Fleurs 64.8 64.2 25.2 20.8 24.7 20.3 27.1 22.1 25.0 20.6 MMS-1B-All 55.2 53.1 21.7 18.1 21.4 17.9 24.9 21.0 21.7 18.2 Whisper-v3-L 78.3 76.3 46.0 42.0 45.5 41.3 39.4 38.4 44.7 40.8 SeamlessM4Tv2-L 87.4 88.5 82.3 79.3 82.2 79.1 55.7 52.9 79.0 75.8 CS Roberta-L-Wh-EN 90.4 91.7 75.8 72.7 73.6 70.8 52.0 48.7 71.5 68.4 Roberta-L-S4T-EN 91.7 90.8 86.2 84.5 85.9 84.2 56.3 55.2 82.3 80.6 LLM2Vec-L-Wh-EN 91.1 92.3 76.6 75.3 74.5 73.4 53.0 53.2 72.4 71.1 LLM2Vec-L-S4T-EN 93.0 91.3 87.1 85.2 86.9 84.9 58.1 56.6 83.4 81.4 NLLB-LLM2Vec-L-Wh 92.5 91.0 77.5 73.9 75.4 72.2 54.4 53.2 73.7 70.4 NLLB-LLM2Vec-L-S4T 94.0 92.7 85.1 84.9 85.3 84.9 62.3 60.7 82.3 81.9 Text Roberta-L-S4T-EN 92.3 87.4 89.1 55.3 85.2 LLM2Vec-S4T-EN 91.0 86.4 87.9 55.6 84.1 NLLB-LLM2Vec 92.3 88.1 87.9 80.7 87.2 Belebele-Fleurs EN (1) Whisper (79) SeamlessM4T (84) Unsupported (4) Not-EN (91) CS Roberta-L-Wh-EN 82.6 81.3 57.5 56.2 56.4 55.2 46.2 44.7 55.5 54.3 Roberta-L-S4T-EN 82.0 79.4 67.5 66.4 67.0 65.8 47.2 46.2 65.2 64.1 LLM2Vec-Wh-EN 94.9 94.4 74.6 73.5 73.6 72.6 63.5 62.9 72.7 71.8 LLM2Vec-S4T-EN 95.5 94.4 84.0 83.1 83.5 82.5 64.7 63.6 81.8 80.9 NLLB-LLM2Vec-Wh 94.8 93.5 58.0 56.2 56.8 55.0 40.9 41.0 55.5 53.8 NLLB-LLM2Vec-S4T 94.9 93.5 61.3 60.2 60.8 59.6 43.9 41.1 59.2 57.8 Text Roberta-L-S4T-EN 83.3 72.1 72.6 48.3 70.5 LLM2Vec-S4T-EN 95.3 87.8 88.3 66.0 86.3 NLLB-LLM2Vec 95.1 65.6 64.6 55.2 63.8
 Table 3 summarizes the main results for both tasks by approach across our xlt setups. We dissect the results along several axes of analysis.
-English. The first column presents the in-language (i.e., no-xlt) English performance by task and utterance quality for our models.
-CS.
+English. The first column presents the in-language English performance by task, modality (Text vs. mSE and CS), and utterance quality by model.
+Text & CS.
 English ASR performance is strong for both transcription models (cf.
 Figure 1), and CS effectively leverages the NLU capabilities of LLMs, thanks to their extensive pre-training on both tasks.
+CS are even on par with models trained on gold text, underscoring the quality of ASR on English.
 As a result, CS outperforms all mSE models on SIB-Fleurs.
 On Belebele-Fleurs, both CS models backed by LLM2Vec and NLLB-LLM2Vec perform comparably.
-In sum, the transcription model appears to have little to no impact on English performance.
+In sum, the transcription model has no clear impact on English performance.
 mSE.
 The English results on SIB-Fleurs suggest that the SLU capabilities of mSE models are strongly shaped by their pre-training curriculum.
-MMS-1b, which is pre-trained solely in a self-supervised manner, underperforms other mse on SIB-Fleurs (44.6%).
-Post-hoc fine-tuning of MMS-1b on ASR, either on Fleurs (MMS-1B-Fleurs) or Fleurs combined with additional data (MMS-1b-All), results in significant improvements on SIB-Fleurs of +10-20 percentage points (pp).
-Notably, MMS-1B-Fleurs outperforms MMS-1b-All (+9.6 pp), suggesting that the broader domain mixture in the MMS-1b-All training set negatively affects this downstream task.
+mHubert (41.1 %) and MMS-1b (44.6%), which are pre-trained solely in a self-supervised manner, underperform other mSE on SIB-Fleurs.
+Fine-tuning MMS-1b on ASR, either on Fleurs (MMS-1B-Fleurs) or Fleurs combined with additional data (MMS-1b-All), results in significant improvements on SIB-Fleurs of +10-20 percentage points (pp).
+Notably, MMS-1B-Fleurs outperforms MMS-1b-All (+9.6 pp), suggesting that the broader domain mixture in the MMS-1b-All training set negatively affects this task.
 Whisper-v3-Large surpasses all MMS-1B variants (78.3%).
@@
 -3 pp).
-The joint pre-training on multilingual ASR as well as multilingual and cross-modal MT with text-to-speech knowledge distillation significantly enriches the semantics in speech representations of SeamlessM4Tv2-Large.
-The results particularly highlight that (i) multilingual ASR, (ii) S2ETT pre-training, and (iii) text-to-speech distillation are crucial training objectives for enabling speech encoders to acquire strong SLU capabilities.
+The joint pre-training on multilingual ASR as well as multilingual and cross-modal MT with text-to-speech knowledge distillation (KD) significantly enriches the semantics in speech representations of SeamlessM4Tv2-Large.
+The results particularly highlight that (i) multilingual ASR, (ii) S2ETT pre-training, and (iii) text-to-speech KD are crucial training objectives for enabling speech encoders to acquire strong SLU capabilities.
 Utterance Quality.
-The quality of English utterances does not seem to affect downstream performance for either mse or cs across both tasks.
-This robustness can be attributed to the exhaustive training of all backbone models on English ASR, supported by the large-scale availability of diverse English speech data.
+The quality of English utterances does not affect performance for either mSE or CS across both tasks.
+This can be attributed to the exhaustive training of all models on English ASR due to the scale of English speech training data.
+This is why CS are also on par with Text.
 zs-xlt.
-The right-hand side of Table 1 reports zs-xlt performance in various setups.
+The right-hand side of Table 3 reports zs-xlt performance.
 For each task, we group the languages into (i) languages supported by Whisper-v3, (ii) languages supported by SeamlessM4Tv2, and (iii) languages unsupported by either model.
-The sets of languages of Fleurs that Whisper-v3 and SeamlessM4Tv2 support overlap in 81 languages.131313For unsupported languages, we hand-select the closest supported language to transcribe into.
-CS.
+Whisper-v3 and SeamlessM4Tv2 overlap in their support for 81 Fleurs languages.121212For unsupported languages, we hand-select the closest supported language to transcribe into.
+Text & CS.
 Across both tasks and all language groups, models trained on transcriptions of SeamlessM4Tv2-Large consistently outperform models fine-tuned on transcriptions of Whisper-v3-Large.
@@
 ‘SeamlessM4Tv2’ and ‘unsupported’ language groups in Table 3).
-For SIB-Fleurs, the cs of SeamlessM4Tv2-Large and Roberta-Large deteriorates only slightly in xlt performance to all 101 target language, on average, relative to English (ca.
--9.6 pp).
-Most of this xlt gap comes from the 9 target languages that SeamlessM4Tv2 does not support (-25.6 pp on average).
-In contrast, pairing Roberta-Large with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (-10.4 pp) and a much larger deficit on languages neither model supports (-44.6 pp).
+Nevertheless, while CS are mostly competitive on SIB-Fleurs, they trail models evaluated on ground-truth paragraphs more significantly on Belebele-Fleurs (ca.
+-5 pp).
+We presume that transcription and S2ETT errors propagate more severely in multiple-choice QA.
+For SIB-Fleurs, all CS with SeamlessM4Tv2-Large deteriorate only slightly in xlt performance to all 101 target language, on average, relative to English (ca.
+-10 pp).
+Most of this xlt gap comes from the 7 target languages that SeamlessM4Tv2 does not support (ca.
+-35 pp on average).
+In contrast, pairing LMs with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (ca.
+-13 pp) and again a larger deficit on languages neither model supports (ca.
+-32 pp).
+CS are highly competitive to Text (ca.
+-0.7 pp).
+Only NLLB-LLM2Vec evaluated on gold text tremendously outperforms all other models on unsupported languages, since they are supported by NLLB Team et al.
+(2022).
 For Belebele-Fleurs, S2ETT paired with LLM2Vec (i.e., TTest) outperforms zs-xlt cascading on in-language transcriptions and NLLB-LLM2Vec, except for unsupported languages.
-This suggests that S2ETT of both Whisper-v3 and SeamlesseM4Tv2 sufficiently translates the target languages into English for successful NLU.
+This suggests that S2ETT of both Whisper-v3 and SeamlessM4Tv2 sufficiently translates the target languages into English for successful NLU.
 On unsupported languages, however, NLLB-LLM2Vec performs better for likely two reasons.
@@
 Second, in addition to translation, NLLB was pre-trained with denoising autoencoding, making NLLB-LLM2Vec more resistant to noisy inputs than LLM2Vec.
-The findings suggests that SeamlessM4Tv2-Large is a much more robust transcription model for both in-language ASR as well as S2ETT (cf.
+Overall, the findings suggest that SeamlessM4Tv2-Large is more robust for both in-language ASR as well as S2ETT (cf.
 Figure 1).
@@
 +1-3 pp).
-Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English (i.e., in-language) performance to other languages effectively, with the xlt gap ranging from -32.4 pp for supported languages to -42.4 pp for unseen languages.
-In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp) while maintaining reasonable performance on unsupported languages (65.9%).
-These findings indicate that multilingual cross-modal machine translation and multilingual text-to-speech token distillation align and semantically enrich multilingual speech representations to significantly enhance cross-lingual SLU.
+Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English in-language performance to other languages effectively, with the xlt gap ranging from -32.3 pp for supported languages to -38.9 pp for unseen languages.
+In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp) though also deflates in performance on unsupported languages (55.7%).
+Overall, these findings nevertheless indicate that multilingual cross-modal translation and multilingual text-to-speech distillation align and semantically enrich multilingual speech representations to enhance cross-lingual SLU.
 Utterance Quality.
@@
 Table 3), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages.
-We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.141414We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large.
+We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.131313We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large.
 We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model is highly comparable across splits.
-We trust that neither of the models has been trained on the dev and test set of Fleurs.
-We use the implementations for CER and sacreBELU of Huggingface’ evaluate.
+We trust that the models have not been trained on the dev and test set of Fleurs.
+We use Huggingface’ evaluate to compute for CER and sacreBELU.
 We then ‘macro-average’ the metrics over all languages.
@@
 ‘win ratio’).
-Nevertheless, SeamlessM4Tv2-Large is the overall more robust transcription model: in transcription for languages other than English, the CER of Whisper-v3-large is about twice as high as the CER of SeamlessM4Tv2-Large, on average.
+Nevertheless, SeamlessM4Tv2-Large is the overall more robust transcription model: in transcription for languages other than English, the CER of Whisper-v3-large is more than twice as high as the CER of SeamlessM4Tv2-Large, on average.
 The full per-language results (cf.
@@
 The more robust ASR and much stronger translation performance of SeamlessM4Tv2-Large results stem from its pre-training.
-SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB-200 translation model Team et al.
+SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB translation model Team et al.
 (2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on wav2vec-BERT 2.0 objective.
@@
 Length Adaptor.
-The sequence length of encoded speech typically far exceeds the sequence length of embedded tokenized text for the same input.
-The mse of SeamlessM4Tv2-Large therefore appends a temporal convolution, called the length adaptor, as its final layer to reduce the resolution of speech frames by a factor of 8 Zhao et al.
-(2022); Seamless Communication et al.
-(2023).151515For architectural details on the length adaptor, we refer to the original paper Communication et al.
+The sequence length of encoded speech typically far exceeds the length of embedded tokenized text for the same input.
+The mSE of SeamlessM4Tv2-Large thus appends a temporal convolution, a ‘length adaptor’, as its final layer to reduce the resolution of speech frames by a factor of 8 and to better align the modalities Seamless Communication et al.
+(2023).141414For architectural details on the length adaptor, we refer to the original paper Communication et al.
 (2023).
-For bimodal input, the length adaptor thereby mitigates the discrepancy in sequence lengths between speech encodings and tokenized text embeddings, intending to better align representations across modalities.
-This raises the question of whether a length adaptor is essential for pooling semantic meaning from speech output tokens.
-To investigate this, we ablate SeamlessM4Tv2-Large’s performance on SIB-Fleurs, comparing results with and without the length adaptor.161616We run a full grid search over learning rates (cf.
-§4) for SeamlessM4Tv2-Large without the length adaptor.
-Table 2 presents the results for both variants.
-The results are mixed.
+This begs the question whether a length adaptor is essential for pooling semantics from speech output tokens.
+We hence compare SeamlessM4Tv2-Large’s performance on SIB-Fleurs with and without the length adaptor.
+Table 4 presents the inconclusive results for both variants.
 For English, performance improves on high-quality utterances but decreases on lower-quality ones.
-In contrast, zs-xlt performance shows a slight but consistent decline, which is more pronounced for lower-quality utterances in supported languages and, irrespective of utterance quality, in unsupported languages.
+In contrast, performance most pronouncedly declines for low-quality utterances in zs-xlt to supported languages (-2.6%).
+The zs-xlt unsupported languages is not affected for high-quality utterances (-0.1%) and only slightly for low-quality ones (-0.9%).
 Three factors may explain this modest gap.
-First, removing the length adaptor reduces model capacity by 47M parameters (-7.4 pp), which were subjected to substantial pre-training.
-Second, the length adaptor, as the final layer of the speech encoder, is most explicitly trained to embed the input into a shared multilingual representation space, attended to by the text decoder at every layer.
-Finally, the length adaptor may improve the robustness of speech encodings by filtering noisy frames through temporal downsampling.
-While a modest gap persists across setups, we however cannot decisively infer based on the results whether appending a length adaptor is, as opposed to replacing model capacity with other layers of equal parameter size during pre-training, crucial for improved multilingual SLU.
-We conclude that the pre-training regime is more vital for multilingual SLU abilities of mse than nuanced architectural design choices.
+First, removing the length adaptor reduces model size by 47M parameters (-7.4 pp), which were part of substantial pre-training.
+Second, the length adaptor, as the final layer of the mSE, is most explicitly trained to embed speech into a shared multilingual space, attended to by the text decoder at every layer.
+Finally, the length adaptor might filter noisy frames through temporal downsampling to improve the robustness of speech encodings.
+While a modest gap persists across setups, we cannot decisively infer whether appending a length adaptor is, as opposed to replacing model capacity with other layers of equal parameter size during pre-training, crucial for improved multilingual SLU.
+We conclude that the pre-training regime is more vital for multilingual SLU capabilities of mSE than nuanced architectural design choices.
 ZS-XLT EN (1) S4T (89) N/A (12) AVG (101) S4Tv2-L Size B W B W B W B W Incl.
-LA 635M 87.4 88.5 82.2 79.1 55.0 51.4 79.0 75.8 Excl.
-LA 588M 88.5 86.4 81.0 77.6 52.7 49.1 77.8 74.4 47M +1.1 -2.1 -1.2 -2.5 -2.3 -2.3 -1.2 -1.4
+LA 635M 87.4 88.5 82.2 79.1 55.7 52.9 79.0 75.8 Excl.
+LA 588M 88.5 86.4 80.9 77.5 55.6 52.0 77.8 74.4 47M +1.1 -2.1 -1.3 -2.6 -0.1 -0.9 -1.2 -1.4
 6 Conclusion
@@
 We hope that our findings inspire future work towards developing more efficient multilingual speech encoders that are jointly pre-trained for both multilingual ASR and SLU to close the performance gap between end-to-end speech classification and cascaded approaches.
+Limitations
+Our work has two primary limitations.
+First, the selection of models in CS can be further expanded.
+In particular, additional evaluations on CS that combine transcription models with prompting 70B LLMs would provide valuable insights given the growing interest in instruction fine-tuned LLMs.
+Moreover, assessing speech-LLMs on both tasks would further illuminate multimodal alignment in massively multilingual scenarios.
+Notably, LLM2Vec on Llama 3.1 8B, which underwent extensive task fine-tuning on 63K instances, represents an exceptionally strong language model backbone for CS.
+Compared to concurrent work, our models would have even outperformed prompting 70B LLMs on speech transcriptions in Belebele Costa-jussà et al.
+(2024).
+Additionally, current speech LLMs rely on Whisper, which is inherently limited to 30-second speech segments.
+This limitation hinders their application in Belebele-Fleurs, where the average speech length is 47 seconds (cf.
+Table 2).
+Our compute infrastructure was unable to support prompting such large LLMs on long contexts, exceeding half a minute of speech combined with corresponding questions and answers.
+We therefore leave these evaluations for future work.
+Furthermore, although our findings indicate a clear link between robust multilingual SLU and strong multilingual ASR, future research must more carefully isolate the mechanisms underlying this interaction.
+Such investigations should clarify the conditions (e.g., training objectives, data and language mixtures) under which SLU most effectively contributes to more robust multilingual ASR.
+Nonetheless, our analysis strongly suggests that multilingual SLU and multilingual ASR are mutually beneficial.
+Acknowledgments
+We use AI assistance, specifically chatGPT o3-mini, to refine both the writing and the tables in our manuscript.
 Appendix A Appendix
-A.1 Code
-The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
+A.1 Further Experimental Details
+Compute infrastructure.
+All experiments run on a single Nvidia L40S 48GB or A100 80GB, respectively.
+We estimate that the total compute budget accumulates to about 2K GPU hours.
+Training runs on SIB-Fleurs require about 30-45 minutes, while fine-tuning LLMs for Belebele require roughly 9 hours per run.
+Evaluation, per checkpoint, across all languages supported by the corresponding benchmark, requires about 1 hour due to the comprehensive setups (e.g., original text and two types of transcriptions for speech recordings).
+Code. The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
 A.2 SIB-Fleurs
-Language Train Validation Test afr_Latn 406 86 95 amh_Ethi 752 54 149 arb_Arab 579 64 133 asm_Beng 730 71 176 ast_Latn 701 69 177 azj_Latn 712 71 174 bel_Cyrl 690 71 177 bul_Cyrl 749 70 176 ben_Beng 742 71 176 bos_Latn 746 71 177 cat_Latn 683 71 177 ceb_Latn 741 61 149 ckb_Arab 738 70 176 zho_Hans 751 71 176 ces_Latn 732 68 172 cym_Latn 739 71 177 dan_Latn 696 70 177 deu_Latn 736 69 175 ell_Grek 750 67 168 eng_Latn 738 71 177 spa_Latn 676 71 177 est_Latn 700 71 176 pes_Arab 692 66 165 fin_Latn 735 71 175 tgl_Latn 604 71 176 fra_Latn 753 65 164 gle_Latn 731 71 176 glg_Latn 660 71 174 guj_Gujr 752 71 177 hau_Latn 753 70 166 heb_Hebr 754 70 175 hin_Deva 653 60 132 hrv_Latn 756 71 176 hun_Latn 750 71 177 hye_Armn 741 71 177 ind_Latn 728 69 167 ibo_Latn 737 71 177 isl_Latn 381 18 23 ita_Latn 743 69 175 jpn_Jpan 662 62 164 jav_Latn 740 67 171 kat_Geor 557 69 177 kam_Latn 752 69 179 kea_Latn 725 71 175 kaz_Cyrl 749 70 176 khm_Khmr 588 69 168 kan_Knda 660 70 174 kor_Hang 669 61 141 kir_Cyrl 729 71 177 ltz_Latn 703 71 176 lug_Latn 691 70 173
-Language Train Validation Test lin_Latn 755 59 139 lao_Laoo 591 54 132 lit_Latn 730 71 178 luo_Latn 698 39 98 lvs_Latn 634 69 174 mri_Latn 749 71 176 mkd_Cyrl 680 71 177 mal_Mlym 723 68 174 khk_Cyrl 743 71 177 mar_Deva 749 71 177 zsm_Latn 713 67 171 mlt_Latn 731 71 176 mya_Mymr 746 71 175 nob_Latn 723 51 127 npi_Deva 754 70 175 nld_Latn 729 58 123 nso_Latn 633 70 169 nya_Latn 720 68 169 oci_Latn 756 71 177 ory_Orya 442 71 168 pan_Guru 580 56 143 pol_Latn 723 68 165 pbt_Arab 701 55 144 por_Latn 728 70 177 ron_Latn 734 69 177 rus_Cyrl 733 71 173 snd_Arab 749 71 177 slk_Latn 628 71 169 slv_Latn 704 71 174 sna_Latn 689 71 176 som_Latn 746 70 177 srp_Cyrl 730 63 164 swe_Latn 686 71 168 swh_Latn 745 65 154 tam_Taml 693 71 169 tel_Telu 658 66 153 tgk_Cyrl 680 69 163 tha_Thai 710 71 176 tur_Latn 692 67 164 ukr_Cyrl 732 67 164 umb_Latn 473 39 108 urd_Arab 636 65 120 uzn_Latn 734 69 175 vie_Latn 737 70 176 wol_Latn 643 52 123 xho_Latn 756 71 177 yor_Latn 686 71 172 zho_Hant 624 70 172 zul_Latn 739 69 175 fuv_Latn 752 68 166 gaz_Latn 574 6 17
+Language Code Language Name Train Validation Test afr_Latn Afrikaans 406 86 95 amh_Ethi Amharic 752 54 149 arb_Arab Modern Standard 579 64 133 asm_Beng Assamese 730 71 176 ast_Latn Asturian 701 69 177 azj_Latn North Azerbaijani 712 71 174 bel_Cyrl Belarusian 690 71 177 ben_Beng Bengali 742 71 176 bos_Latn Bosnian 746 71 177 bul_Cyrl Bulgarian 749 70 176 cat_Latn Catalan 683 71 177 ceb_Latn Cebuano 741 61 149 ces_Latn Czech 732 68 172 ckb_Arab Central Kurdish 738 70 176 cym_Latn Welsh 739 71 177 dan_Latn Danish 696 70 177 deu_Latn German 736 69 175 ell_Grek Greek 750 67 168 eng_Latn English 738 71 177 est_Latn Estonian 700 71 176 fin_Latn Finnish 735 71 175 fra_Latn French 753 65 164 fuv_Latn Nigerian Fulfulde 752 68 166 gaz_Latn West Central Oromo 574 6 17 gle_Latn Irish 731 71 176 glg_Latn Galician 660 71 174 guj_Gujr Gujarati 752 71 177 hau_Latn Hausa 753 70 166 heb_Hebr Hebrew 754 70 175 hin_Deva Hindi 653 60 132 hrv_Latn Croatian 756 71 176 hun_Latn Hungarian 750 71 177 hye_Armn Armenian 741 71 177 ibo_Latn Igbo 737 71 177 ind_Latn Indonesian 728 69 167 isl_Latn Icelandic 381 18 23 ita_Latn Italian 743 69 175 jav_Latn Javanese 740 67 171 jpn_Jpan Japanese 662 62 164 kam_Latn Kamba 752 69 179 kan_Knda Kannada 660 70 174 kat_Geor Georgian 557 69 177 kaz_Cyrl Kazakh 749 70 176 kea_Latn Kabuverdianu 725 71 175 khk_Cyrl Halh 743 71 177 khm_Khmr Khmer 588 69 168 kir_Cyrl Kyrgyz 729 71 177 kor_Hang Korean 669 61 141 lao_Laoo Lao 591 54 132 lin_Latn Lingala 755 59 139 lit_Latn Lithuanian 730 71 178
+Language Code Language Name Train Validation Test ltz_Latn Luxembourgish 703 71 176 lug_Latn Ganda 691 70 173 luo_Latn Luo 698 39 98 lvs_Latn Standard Latvian 634 69 174 mal_Mlym Malayalam 723 68 174 mar_Deva Marathi 749 71 177 mkd_Cyrl Macedonian 680 71 177 mlt_Latn Maltese 731 71 176 mri_Latn Maori 749 71 176 mya_Mymr Burmese 746 71 175 nld_Latn Dutch 729 58 123 nob_Latn Norwegian Bokmål 723 51 127 npi_Deva Nepali 754 70 175 nso_Latn Northern Sotho 633 70 169 nya_Latn Nyanja 720 68 169 oci_Latn Occitan 756 71 177 ory_Orya Odia 442 71 168 pan_Guru Eastern Panjabi 580 56 143 pbt_Arab Southern Pashto 701 55 144 pes_Arab Western Persian 692 66 165 pol_Latn Polish 723 68 165 por_Latn Portuguese 728 70 177 ron_Latn Romanian 734 69 177 rus_Cyrl Russian 733 71 173 slk_Latn Slovak 628 71 169 slv_Latn Slovenian 704 71 174 sna_Latn Shona 689 71 176 snd_Arab Sindhi 749 71 177 som_Latn Somali 746 70 177 spa_Latn Spanish 676 71 177 srp_Cyrl Serbian 730 63 164 swe_Latn Swedish 686 71 168 swh_Latn Swahili 745 65 154 tam_Taml Tamil 693 71 169 tel_Telu Telugu 658 66 153 tgk_Cyrl Tajik 680 69 163 tgl_Latn Tagalog 604 71 176 tha_Thai Thai 710 71 176 tur_Latn Turkish 692 67 164 ukr_Cyrl Ukrainian 732 67 164 umb_Latn Umbundu 473 39 108 urd_Arab Urdu 636 65 120 uzn_Latn Northern Uzbek 734 69 175 vie_Latn Vietnamese 737 70 176 wol_Latn Wolof 643 52 123 xho_Latn Xhosa 756 71 177 yor_Latn Yoruba 686 71 172 zho_Hans Chinese 751 71 176 zho_Hant Chinese 624 70 172 zsm_Latn Standard Malay 713 67 171 zul_Latn Zulu 739 69 175
 A.3 Belebele-Fleurs
-Language Samples afr_Latn 309 amh_Ethi 782 arb_Arab 387 asm_Beng 824 azj_Latn 759 ben_Beng 855 bul_Cyrl 873 cat_Latn 652 ceb_Latn 783 ces_Latn 802 ckb_Arab 842 dan_Latn 696 deu_Latn 804 ell_Grek 837 eng_Latn 844 est_Latn 736 fin_Latn 826 fra_Latn 839 fuv_Latn 848 gaz_Latn 25 guj_Gujr 880 hau_Latn 838 heb_Hebr 878 hin_Deva 515 hrv_Latn 896 hun_Latn 879 hye_Armn 861 ibo_Latn 838 ind_Latn 783 isl_Latn 81 ita_Latn 851 jav_Latn 835 jpn_Jpan 590 kan_Knda 606 kat_Geor 372 kaz_Cyrl 870 kea_Latn 770 khk_Cyrl 869 khm_Khmr 439 kir_Cyrl 811 kor_Hang 535 lao_Laoo 346 lin_Latn 778 lit_Latn 834 lug_Latn 703 luo_Latn 512
-Language Samples lvs_Latn 555 mal_Mlym 809 mar_Deva 869 mkd_Cyrl 667 mlt_Latn 816 mri_Latn 877 mya_Mymr 864 nld_Latn 674 nob_Latn 635 npi_Deva 876 nso_Latn 569 nya_Latn 752 ory_Orya 220 pan_Guru 396 pbt_Arab 628 pes_Arab 673 pol_Latn 765 por_Latn 791 ron_Latn 815 rus_Cyrl 819 slk_Latn 513 slv_Latn 724 sna_Latn 735 snd_Arab 878 som_Latn 874 spa_Latn 659 srp_Cyrl 766 swe_Latn 681 swh_Latn 780 tam_Taml 714 tel_Telu 567 tgk_Cyrl 632 tgl_Latn 505 tha_Thai 745 tur_Latn 706 ukr_Cyrl 773 urd_Arab 482 uzn_Latn 812 vie_Latn 847 wol_Latn 495 xho_Latn 900 yor_Latn 652 zho_Hans 888 zho_Hant 527 zsm_Latn 749 zul_Latn 838
+Language Code Language Name Samples afr_Latn Afrikaans 309 amh_Ethi Amharic 782 arb_Arab Modern Standard 387 asm_Beng Assamese 824 azj_Latn North Azerbaijani 759 ben_Beng Bengali 855 bul_Cyrl Bulgarian 873 cat_Latn Catalan 652 ceb_Latn Cebuano 783 ces_Latn Czech 802 ckb_Arab Central Kurdish 842 dan_Latn Danish 696 deu_Latn German 804 ell_Grek Greek 837 eng_Latn English 844 est_Latn Estonian 736 fin_Latn Finnish 826 fra_Latn French 839 fuv_Latn Nigerian Fulfulde 848 gaz_Latn West Central Oromo 25 guj_Gujr Gujarati 880 hau_Latn Hausa 838 heb_Hebr Hebrew 878 hin_Deva Hindi 515 hrv_Latn Croatian 896 hun_Latn Hungarian 879 hye_Armn Armenian 861 ibo_Latn Igbo 838 ind_Latn Indonesian 783 isl_Latn Icelandic 81 ita_Latn Italian 851 jav_Latn Javanese 835 jpn_Jpan Japanese 590 kan_Knda Kannada 606 kat_Geor Georgian 372 kaz_Cyrl Kazakh 870 kea_Latn Kabuverdianu 770 khk_Cyrl Halh 869 khm_Khmr Khmer 439 kir_Cyrl Kyrgyz 811 kor_Hang Korean 535 lao_Laoo Lao 346 lin_Latn Lingala 778 lit_Latn Lithuanian 834 lug_Latn Ganda 703 luo_Latn Luo 512 lvs_Latn Standard Latvian 555 mal_Mlym Malayalam 809
+Language Name Samples mar_Deva Marathi 869 mkd_Cyrl Macedonian 667 mlt_Latn Maltese 816 mri_Latn Maori 877 mya_Mymr Burmese 864 nld_Latn Dutch 674 nob_Latn Norwegian Bokmål 635 npi_Deva Nepali 876 nso_Latn Northern Sotho 569 nya_Latn Nyanja 752 ory_Orya Odia 220 pan_Guru Eastern Panjabi 396 pbt_Arab Southern Pashto 628 pes_Arab Western Persian 673 pol_Latn Polish 765 por_Latn Portuguese 791 ron_Latn Romanian 815 rus_Cyrl Russian 819 slk_Latn Slovak 513 slv_Latn Slovenian 724 sna_Latn Shona 735 snd_Arab Sindhi 878 som_Latn Somali 874 spa_Latn Spanish 659 srp_Cyrl Serbian 766 swe_Latn Swedish 681 swh_Latn Swahili 780 tam_Taml Tamil 714 tel_Telu Telugu 567 tgk_Cyrl Tajik 632 tgl_Latn Tagalog 505 tha_Thai Thai 745 tur_Latn Turkish 706 ukr_Cyrl Ukrainian 773 urd_Arab Urdu 482 uzn_Latn Northern Uzbek 812 vie_Latn Vietnamese 847 wol_Latn Wolof 495 xho_Latn Xhosa 900 yor_Latn Yoruba 652 zho_Hans Chinese 888 zho_Hant Chinese 527 zsm_Latn Standard Malay 749 zul_Latn Zulu 838
 A.4 Analysis of Speech-To-English-Text-Translation Performance
@@
 A.5 Analysis of ASR Performance
-Language Whisper-v3-Large SeamlessM4Tv2-Large afr_Latn 13.1 12.6 amh_Ethi 207.8 23.9 arb_Arab 7.6 8.0 asm_Beng 99.0 15.1 ast_Latn 15.5 19.8 azj_Latn 7.2 8.3 bel_Cyrl 11.9 6.9 ben_Beng 33.9 9.0 bos_Latn 4.6 7.8 bul_Cyrl 5.2 9.6 cat_Latn 3.3 5.3 ceb_Latn 17.4 19.0 ces_Latn 4.2 8.0 ckb_Arab 98.3 13.9 cym_Latn 15.8 14.4 dan_Latn 4.8 10.0 deu_Latn 2.2 6.5 ell_Grek 6.4 10.8 eng_Latn 3.1 6.2 est_Latn 7.3 11.0 fin_Latn 3.2 10.2 fra_Latn 2.8 6.6 fuv_Latn 194.3 57.0 gaz_Latn 35.8 47.8 gle_Latn 85.1 21.4 glg_Latn 4.3 6.5 guj_Gujr 21.3 10.3 hau_Latn 33.8 51.8 heb_Hebr 12.2 14.5 hin_Deva 11.7 10.2 hrv_Latn 4.5 8.5 hun_Latn 5.2 8.5 hye_Armn 18.0 9.1 ibo_Latn 42.9 59.3 ind_Latn 4.0 9.2 isl_Latn 16.0 10.3 ita_Latn 2.1 5.2 jav_Latn 26.3 11.3 jpn_Jpan 6.3 16.4 kam_Latn 45.1 53.9 kan_Knda 20.1 10.9 kat_Geor 19.8 7.1 kaz_Cyrl 8.9 9.7 kea_Latn 35.8 38.1 khk_Cyrl 37.6 14.6 khm_Khmr 144.0 29.8 kir_Cyrl 28.3 8.8 kor_Hang 7.4 10.3 lao_Laoo 109.7 30.3 lin_Latn 22.0 58.7 lit_Latn 8.3 11.9 ltz_Latn 29.5 40.1
-Language Whisper-v3-Large SeamlessM4Tv2-Large lug_Latn 44.9 14.1 luo_Latn 42.7 58.8 lvs_Latn 7.3 8.2 mal_Mlym 104.3 14.2 mar_Deva 24.3 10.7 mkd_Cyrl 5.8 8.6 mlt_Latn 26.0 12.1 mri_Latn 13.4 54.7 mya_Mymr 132.9 22.2 nld_Latn 3.1 6.8 nob_Latn 4.6 9.0 npi_Deva 25.6 74.1 nso_Latn 95.7 66.9 nya_Latn 35.6 12.4 oci_Latn 25.9 33.8 ory_Orya 93.8 11.4 pan_Guru 44.2 10.3 pbt_Arab 36.8 21.6 pes_Arab 8.6 7.7 pol_Latn 2.3 7.9 por_Latn 2.8 7.6 ron_Latn 3.2 6.6 rus_Cyrl 2.6 6.1 slk_Latn 3.9 7.0 slv_Latn 6.0 9.2 sna_Latn 26.7 32.7 snd_Arab 94.5 33.5 som_Latn 34.9 17.9 spa_Latn 2.2 6.3 srp_Cyrl 76.6 7.6 swe_Latn 3.7 9.2 swh_Latn 11.0 8.9 tam_Taml 13.3 11.6 tel_Telu 90.1 13.3 tgk_Cyrl 28.8 8.8 tgl_Latn 5.0 12.0 tha_Thai 9.8 11.4 tur_Latn 6.2 7.5 ukr_Cyrl 3.0 8.9 umb_Latn 156.7 55.5 urd_Arab 9.4 9.6 uzn_Latn 25.0 8.4 vie_Latn 4.7 6.7 wol_Latn 140.6 47.9 xho_Latn 48.1 32.3 yor_Latn 49.3 33.2 zho_Hans 16.2 19.0 zho_Hant 30.6 35.1 zsm_Latn 3.4 8.5 zul_Latn 42.6 18.4
+Language Whisper-v3-Large SeamlessM4Tv2-Large afr_Latn 13.1 12.6 amh_Ethi 207.8 23.9 arb_Arab 7.6 8.0 asm_Beng 99.0 15.1 ast_Latn 15.5 19.8 azj_Latn 7.2 8.3 bel_Cyrl 11.9 6.9 ben_Beng 33.9 9.0 bos_Latn 4.6 7.8 bul_Cyrl 5.2 9.6 cat_Latn 3.3 5.3 ceb_Latn 17.4 19.0 ces_Latn 4.2 8.0 ckb_Arab 98.3 13.9 cym_Latn 15.8 14.4 dan_Latn 4.8 10.0 deu_Latn 2.2 6.5 ell_Grek 6.4 10.8 eng_Latn 3.1 6.2 est_Latn 7.3 11.0 fin_Latn 3.2 10.2 fra_Latn 2.8 6.6 fuv_Latn 194.3 57.0 gaz_Latn 35.8 47.8 gle_Latn 85.1 21.4 glg_Latn 4.3 6.5 guj_Gujr 21.3 10.3 hau_Latn 33.8 51.8 heb_Hebr 12.2 14.5 hin_Deva 11.7 10.2 hrv_Latn 4.5 8.5 hun_Latn 5.2 8.5 hye_Armn 18.0 9.1 ibo_Latn 42.9 59.3 ind_Latn 4.0 9.2 isl_Latn 16.0 10.3 ita_Latn 2.1 5.2 jav_Latn 26.3 11.3 jpn_Jpan 6.3 16.4 kam_Latn 45.1 53.9 kan_Knda 20.1 10.9 kat_Geor 19.8 7.1 kaz_Cyrl 8.9 9.7 kea_Latn 35.8 38.1 khk_Cyrl 37.6 14.6 khm_Khmr 144.0 29.8 kir_Cyrl 28.3 8.8 kor_Hang 7.4 10.3 lao_Laoo 109.7 30.3 lin_Latn 22.0 58.7 lit_Latn 8.3 11.9
+Language Whisper-v3-Large SeamlessM4Tv2-Large ltz_Latn 29.5 40.1 lug_Latn 44.9 14.1 luo_Latn 42.7 58.8 lvs_Latn 7.3 8.2 mal_Mlym 104.3 14.2 mar_Deva 24.3 10.7 mkd_Cyrl 5.8 8.6 mlt_Latn 26.0 12.1 mri_Latn 13.4 54.7 mya_Mymr 132.9 22.2 nld_Latn 3.1 6.8 nob_Latn 4.6 9.0 npi_Deva 25.6 74.1 nso_Latn 95.7 66.9 nya_Latn 35.6 12.4 oci_Latn 25.9 33.8 ory_Orya 93.8 11.4 pan_Guru 44.2 10.3 pbt_Arab 36.8 21.6 pes_Arab 8.6 7.7 pol_Latn 2.3 7.9 por_Latn 2.8 7.6 ron_Latn 3.2 6.6 rus_Cyrl 2.6 6.1 slk_Latn 3.9 7.0 slv_Latn 6.0 9.2 sna_Latn 26.7 32.7 snd_Arab 94.5 33.5 som_Latn 34.9 17.9 spa_Latn 2.2 6.3 srp_Cyrl 76.6 7.6 swe_Latn 3.7 9.2 swh_Latn 11.0 8.9 tam_Taml 13.3 11.6 tel_Telu 90.1 13.3 tgk_Cyrl 28.8 8.8 tgl_Latn 5.0 12.0 tha_Thai 9.8 11.4 tur_Latn 6.2 7.5 ukr_Cyrl 3.0 8.9 umb_Latn 156.7 55.5 urd_Arab 9.4 9.6 uzn_Latn 25.0 8.4 vie_Latn 4.7 6.7 wol_Latn 140.6 47.9 xho_Latn 48.1 32.3 yor_Latn 49.3 33.2 zho_Hans 16.2 19.0 zho_Hant 30.6 35.1 zsm_Latn 3.4 8.5 zul_Latn 42.6 18.4
 A.6 Silent Fleurs Examples

### v3

@@
 Abstract
-Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system, since these languages cannot pair automatic speech recognition (ASR) with language models to benefit from language technology.
+Spoken language understanding (SLU) is indispensable for half of all living languages that lack a formal writing system.
+Unlike for high-resource languages, for these languages, we cannot offload semantic understanding of speech to the cascade of automatic speech recognition (ASR) and text-based large language models (LLMs).
 Even if low-resource languages possess a writing system, ASR for these languages remains unreliable due to limited bimodal speech and text training data.
-Better SLU can strengthen the robustness of massively multilingual ASR by levering language semantics to disambiguate utterances via context or exploiting semantic similarities across languages.
-However, the evaluation of multilingual SLU remains limited to shallow tasks such as intent classification or language identification.
-To address this, we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering through listening comprehension spanning 944 hours of speech across 92 languages.
-We extensively evaluate both end-to-end speech classification models and cascaded systems that combine speech-to-text transcription with subsequent classification by large language models on Fleurs-SLU.
-Our results show that cascaded systems exhibit greater robustness in multilingual SLU tasks, though speech encoders can achieve competitive performance in topical speech classification when appropriately pre-trained.
-We further find a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, highlighting the mutual benefits between acoustic and semantic speech representations.111The datasets are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
-Fleurs-SLU: A Massively Multilingual Benchmark for
-Spoken Language Understanding
-Fabian David Schmidt1, Ivan Vulić2, Goran Glavaš1, David Ifeoluwa Adelani3 1 Center For Artificial Intelligence and Data Science, University of Würzburg, Germany 2 Language Technology Lab, University of Cambridge, United Kingdom 3 Mila, McGill University and Canada CIFAR AI Chair {fabian.schmidt, goran.glavas}@uni-wuerzburg.de iv250@cam.ac.uk david.adelani@mila.quebec
+Nonetheless, the evaluation of multilingual SLU is limited to shallow tasks such as intent classification or language identification.
+This is why we present Fleurs-SLU, a multilingual SLU benchmark that encompasses (i) 692 hours of speech for topical utterance classification in 102 languages and (ii) multiple-choice question answering via listening comprehension spanning 944 hours of speech across 92 languages.
+We extensively evaluate end-to-end speech classification models, cascaded systems that combine speech-to-text transcription with subsequent LLM-based classification, and multimodal speech-LLMs on Fleurs-SLU.
+Our results show that cascaded systems are more robust in multilingual SLU, though well-pretrained speech encoders can perform competitively in topical speech classification.
+Closed-source speech-LLMs match or surpass the performance of cascaded systems.
+We observe a strong correlation between robust multilingual ASR, effective speech-to-text translation, and strong multilingual SLU, indicating mutual benefits between acoustic and semantic speech representations.
+111The datasets & per-language results are available on Huggingface: SIB-Fleurs; Belebele-Fleurs.
 1 Introduction
-Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology Ethnologue (2017).
+Only about half of the world’s living languages possess a formal writing system, underscoring the need for massively multilingual speech technology (Ethnologue, 2017).
 Spoken language understanding (SLU) is hence a critical feature of inclusive technology for these languages, which cannot rely on combining automatic speech recognition (ASR) with language models to offload semantic speech understanding.
 However, the datasets currently available for evaluating SLU in truly relevant languages suffer from significant limitations.
-For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords Gerz et al.
-(2021).
-Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage Keller and Glavaš (2024).
+For instance, the Minds14 benchmark assesses SLU on 14 exclusively high-resource languages for intent classification, a task that often requires only shallow semantic processing, such as detecting specific keywords (Gerz et al., 2021).
+Likewise, while SpeechTaxi focuses on spoken topical classification of Bible verses in 28 diverse languages, SLU of Bible verse arguably does not reflect real-world usage (Keller & Glavaš, 2024).
 Moreover, even when a low-resource language has a writing system, multilingual ASR often struggles to reliably transcribe that language due to the limited availability of bimodal speech and text data.
-Pre-trained on the self-supervised wav2vec-BERT objective Baevski et al.
-(2020), these models embed acoustic features that capture phonetic rather than semantic information Choi et al.
-(2024).
+Pre-trained on the self-supervised wav2vec-BERT objective (Baevski et al., 2020), these models embed acoustic features that capture phonetic rather than semantic information (Choi et al., 2024).
 Conversely, multilingual speech models that excel in semantic encoding tend to be more robust in ASR tasks by, e.g., leveraging context-related cues or cross-lingual similarities such as cognates or similar syntactic structures (cf.
 §5.1).
-This underscores that robust SLU is an instrumental component of genuinely inclusive speech technology.
-To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores-200 (Flores).
-We first filter out silent instances from Fleurs Conneau et al.
-(2022) and then map the remaining data back to Flores Team et al.
-(2022).
-The resulting dataset is then merged with SIB-200 Adelani et al.
-(2024) and Belebele Bandarkar et al.
-(2023) to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
+This underscores that robust SLU is instrumental for genuinely inclusive speech technology.
+To address the shortcomings of multilingual SLU evaluation of prior benchmarks, we compile Fleurs-SLU by realigning datasets derived from Flores (Team et al., 2022), a machine-translation benchmark with parallel sentences for over 200 languages.
+We first filter out silent instances from Fleurs (Conneau et al., 2022) and then map the remaining data back to Flores.
+The resulting dataset is then merged with SIB-200 (Adelani et al., 2024) and Belebele (Bandarkar et al., 2023), that both are based on Flores, to create a topical utterance classification benchmark for 102 languages, and dataset for multiple-choice question answering (QA) on spoken paragraphs for 92 languages, respectively.
 Contributions.
 1) To the best of our knowledge, we are the first to present a multilingual SLU benchmark that spans over 100 languages and enables end-to-end SLU evaluation of multilingual speech encoders (mSE).
-SIB-Fleurs supports semantic utterance classification with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks.
-Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages Costa-jussà et al.
-(2024).
+SIB-Fleurs supports utterance classification for 7 topics with 692 hours of speech spanning 102 languages, over 70 more additional languages than existing benchmarks.
+Belebele-Fleurs provides textual multiple-choice QA from spoken paragraphs, totaling 944 hours of speech across 92 languages and exceeding concurrent work by 18 languages (Costa-jussà et al., 2024).
 2) We extensively benchmark state-of-the-art speech models on SIB-Fleurs and Belebele-Fleurs.
-Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding Keller and Glavaš (2024).
+Like prior work, we observe that Cascaded Systems (CS) remain more robust than mSE, enabling SLU competitive with text-based language understanding (Keller & Glavaš, 2024).
+We additionally benchmark state-of-the-art speech-LLMs on multilingual SLU and find that often perform on par or better than CS.
 We are the first to show that mSE pre-trained on language understanding objectives can yield performance competitive to that of CS on speech classification.
@@
 3) We empirically demonstrate that strong multilingual SLU coincides with much more robust multilingual ASR and higher-quality speech-to-English-text translation (S2ETT).
-This suggests that the pre-training of multilingual speech models should take SLU capabilities into account to make massively multilingual ASR more robust.
+This suggests that the pre-training of multilingual speech models may take SLU into account to make multilingual ASR more robust.
 2 Related Work
@@
 Modern speech representation models are pre-trained with self-supervised objectives that are optionally followed-up by ASR and speech-to-text translation (S2TT) training.
-Among these models, mHubert employs the self-supervised wav2vec-BERT 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features Boito et al.
-(2024).
-MMS-1b is also pretrained with wav2vec-BERT 2.0 on an large corpus of 491K hours of speech spanning 1,406 languages Pratap et al.
-(2024).
-Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio Radford et al.
-(2022).
-SeamlessM4Tv2-Large is a multilingual multimodal translation model Seamless Communication et al.
-(2023).
-It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with wav2vec-BERT 2.0 on 4.5M hours of audio.
+Among these models, mHubert employs the self-supervised wav2vec 2.0 objective on 90K hours of speech across 147 languages by predicting pseudo-labels derived from clustering raw speech features (Boito et al., 2024).
+MMS-1b is also pretrained with wav2vec 2.0 on a large corpus of 491K hours of speech spanning 1,406 languages (Pratap et al., 2024).
+Whisper-v3 is a Transformer encoder-decoder that has been multi-task pre-trained on multilingual ASR, speech-to-English-text-translation (S2ETT), spoken language identification, and Voice activity detection on 680k hours of audio (Radford et al., 2022).
+SeamlessM4Tv2-Large is a multilingual multimodal translation model (Seamless Communication et al., 2023).
+It combines a text-to-text translation (T2TT) model pre-trained on 105 languages and a Conformer speech encoder pre-trained with w2v-BERT 2.0 on 4.5M hours of audio.
 The model is then trained on T2TT, S2TT, ASR, and knowledge distillation (KD) objectives.
@@
 Multilingual exceptions are otherwise limited.
-The Minds14 dataset for intent classification only includes 14 high-resource languages Gerz et al.
-(2021).
-SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains Keller and Glavaš (2024).
+The Minds14 dataset for intent classification only includes 14 high-resource languages (Gerz et al., 2021).
+SpeechTaxi offers spoken topical classification for Bible verses in 28 diverse languages, which however do not adequately represent real-world domains (Keller & Glavaš, 2024).
 Concurrent Work.
@@
 (2024) concurrently released the multilingual SLU benchmark 2M-Belebele for 74 languages.
-In this dataset, the authors first extend Flores by ca.
-20% by incorporating human recordings for both sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele.
+In this dataset, the authors first extend Fleurs by approx.
+20% by incorporating human recordings for sentences that are part of Flores but were missing in Fleurs, as well as the questions and answers from Belebele.
 Costa-jussà et al.
-(2024) find that a CS-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by ca.
+(2024) find that a CS-based approach on transcribing speech with Whisper-v3-Large and subsequently prompting Llama 3 70B with the transcription trails prompting the LLM on clean text by approx.
 8 percentage points (pp).
 3 Fleurs-SLU
-We postprocess datasets that are based on Flores to create ‘Fleurs-SLU’, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs.
+We create Fleurs-SLU, a massively multilingual SLU benchmark for speech classification and multiple-choice QA from spoken paragraphs, from datasets that are based on Flores.
 3.1 Core Datasets
-Flores.
-The authors professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation Team et al.
-(2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets.
+Flores consists of professionally translated 3,001 sentences of English Wikipedia paragraphs to evaluate machine translation (Team et al., 2022).222Flores comprises 3,001 sentences divided into dev (997 sentences), devtest (1,012 sentences), and test (992 sentences) sets.
 The authors did not release the test set.
-Fleurs.
-Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores Conneau et al.
-(2022).333Note that almost all languages part of Fleurs however are missing a few hundred sentences from Flores.
-Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech translation in all language directions.
-SIB-200.
-Adelani et al.
-(2024) refined the topical metadata annotations of sentences in the dev and devtest splits of Flores into 7 categories.444The categories are ”science/technology”, ”travel”, ”politics”, ”sports”, ”health”, ”entertainment”, ”geography”.
-The resulting SIB-200 is a natural language understanding benchmark for 205 language variants.
-Belebele.
-Belebele is a multiple-choice reading comprehension benchmark for 122 languages Bandarkar et al.
-(2024).
+Fleurs comprises 2.3 spoken utterances, on average, per sentence from the dev and devtest splits of 102 languages in Flores (Conneau et al., 2022).333Almost all languages part of Fleurs however are missing a few hundred sentences from Flores.
+Fleurs is used to evaluate multilingual ASR, LID, as well as speech-to-text and text-to-speech tanslation in all language directions.
+SIB-200 refined the topical metadata annotations of sentences in the dev and devtest splits of Flores into 7 categories (Adelani et al., 2024).444”science/technology”, ”travel”, ”politics”, ”sports”, ”health”, ”entertainment”, ”geography”.
+The resulting SIB-200 is a topical classification benchmark for 205 language variants.
+Belebele is a multiple-choice reading comprehension benchmark for 122 languages (Bandarkar et al., 2024).
 The authors reconstruct paragraphs from Flores sentences.
-They generate 1–2 questions per English paragraph, which are professionally translated into 121 languages.
-Belebele comprises 900 questions spanning 488 passages.
+They generate 1-2 questions per English paragraph, which are professionally translated into 121 languages.
+Belebele comprises 900 questions that span across 488 passages.
 3.2 Benchmark Creation
@@
 We begin by removing silent instances from Fleurs.555Counts of removed examples are listed in Appendix A.6.
-We first normalize the loudness to an RMS target level of .
-We next apply voice activity detection using Silero-VAD Team (2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect.
+We first normalize the loudness to a target RMS level of dB.
+We next apply voice activity detection using Silero-VAD (Team, 2024).666https://github.com/snakers4/silero-vad Samples are deemed silent if speech is detected in less than of their duration.777The Silero-VAD pipeline also frequently removes inaudibly noisy samples as a side effect.
 Lastly, we verify our approach on 50 randomly sampled predicted silent instances.
-We find only one borderline misclassified example, which is very noisy but still comprehensible.
+We find only one borderline misclassified example, which is noisy but comprehensible.
 We then conservatively merge Fleurs and Flores by matching instances first on exact string match and then by Levenshtein distance of 3 on normalized strings.888Normalized strings remove characters that are Unicode punctuation codepoints.
 2a) SIB-Fleurs.
-For each language, we pool instances from the training, validation, and test splits of SIB-200 Adelani et al.
-(2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before.
+For each language, we pool instances from the training, validation, and test splits of SIB-200 (Adelani et al., 2024) and align the data with our merged Fleurs-Flores dataset with the same string alignment procedure as before.
 The data is then regrouped into the training, validation, and test splits of the original Fleurs dataset.
@@
 Appendix A.2 provides a list of samples by split per language.
-| Task | Utterance |
+Utterance | |
 Classification | |
@@
 Test | 163 (174) |
+Multiple-Choice QA | |
+Spoken Paragraphs | |
+Languages | 92 |
+Answer Choices | 4 |
+Questions per paragraph | 1.8 (1.8) |
+Sentences per paragraph | 3.6 (4.0) |
+Utterances per sentence | 2.0 (2.0) |
+Duration per utterance (s) | 12.6 (12.0) |
+Duration per paragraph (s) | 47.0 (43.4) |
+Total audio (hr) | 944 |
+Samples (Paragraphs) | 709 (771) |
 2b) Belebele-Fleurs.
@@
 We verify our reconstructed paragraphs against the original Belebele by ensuring that the Levenshtein distance for strings with removed punctuation is negligibly small (less than 3 characters).
-Table 2 provides a detailed summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
-Task Multiple-Choice QA Spoken Paragraphs Languages 92 Answer Choices 4 Questions per paragraph 1.8 (1.8) Sentences per paragraph 3.6 (4.0) Utterances per sentence 2.0 (2.0) Duration per utterance (s) 12.6 (12.0) Duration per paragraph (s) 47.0 (43.4) Total audio (hr) 944 Samples (Paragraphs) 709 (771)
+Table 2 provides a summary statistics on both the paragraph- and the sentence-level for Belebele-Fleurs, while Appendix A.3 lists the samples per language.
 4 Experimental Setup
@@
 Belebele-Fleurs.
-We train and validate models on the English training and dev sets of Bandarkar et al.
-(2024), respectively.
+We train and validate CS models on the English training and dev sets of Belebele (Bandarkar et al., 2024), respectively.
 We jointly embed the paragraph, question, and choices with text encoders.
@@
 Zero-shot cross-lingual transfer (zs-xlt) and Translate-Test (TTest) allow us to evaluate xlt without requiring additional annotation for any target language.
-In zs-xlt, we first train a multilingual model on source-language data and then directly run inference on target-language instances.
-In TTest, the model is first fine-tuned on labeled source-language data.
+In zs-xlt, we first train a multilingual model on the English source-language data (cf.
+§4.1) and then directly run inference on the target-language test instances.
+In TTest, the model is also first fine-tuned on labeled English source-language data.
 At test time, the target-language examples are translated to the source language prior to inference, which enables xlt with monolingual LLMs.
+We additionally evaluate zero-shot prompting of speech-LLMs.
+Here, we provide the context of the task in English and all relevant input in-language in the respective modality, i.e.
+text or utterances, to the speech-LLMs (cf.
+§A.1).
 Speech Classification.
@@
 §2).
-We include MMS-1b without fine-tuning (‘MMS-1b’), with ASR fine-tuning on Fleurs (‘MMS-1b-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1b-all’).
+We include MMS-1B without fine-tuning (‘MMS-1B’), with ASR fine-tuning on Fleurs (‘MMS-1B-Fleurs’), and with ASR fine-tuning on multilingual datasets (‘MMS-1B-all’), allowing us to analyze the impact of ASR fine-tuning on cross-lingual SLU.
 For ASR fine-tuning, Pratap et al.
@@
 For task fine-tuning, we freeze the adapters to facilitate zs-xlt.
-We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large.
+We further evaluate mHubert and the speech encoders of both Whisper-v3-Large and SeamlessM4Tv2-Large (cf.
+§2).
 Cascading & Text.
-Cascaded systems (CS) perform xlt in two steps.
-CS first transcribe speech into text with an ASR model.
-The transcription is then fed into an text encoder paired with a classification head.
-We transcribe speech into the target language and into English with SeamlessM4Tv2-Large and Whisper-v3-Large.
-Note that speech-to-text-translation into English (S2ETT) with CS is equivalent to TTest as merely monolingual LLMs are needed for xlt.
-We further evaluate on the original text of SIB and Bebelebe, which we translate to English with SeamlessM4Tv2-Large.101010We use the gold Q&A for the respective language (i.e., English for ).
-Otherwise, we would need to interleave MT models, e.g., S2ETT of paragraphs with Whisper, but translating textual Q&A with SeamlessM4Tv2.
-We decode greedily in transcription, since more complex decoding strategies do not improve xlt performance Ebing and Glavaš (2024).
-We run our text-based and CS experiments on Roberta-Large Liu et al.
-(2019) as well as LLM2Vec and NLLB-LLM2Vec that are sequence encoders based on Llama 3 8B AI@Meta (2024).
-LLM2Vec BehnamGhader et al.
-(2024) enables bidirectional attention and continues training on both masked next-token prediction and SimCSE Gao et al.
-(2021).
-NLLB-LLM2Vec fuses LLM2Vec with the encoder of the NLLB translation model to extend its NLU capabilities to 200 languages Schmidt et al.
-(2024).
-4.3 Training Setup
-Training Details.
-We train all models with AdamW (Loshchilov and Hutter, 2019), weight decay of , and with 10% linear warm-up and followed by linear decay on an effective batch size of 32.
+Cascaded systems (CS) perform xlt in two steps: (i) an ASR model transcribes speech into text, and (ii) a text encoder processes the transcription via a classification head.
+We consider two transcription targets: the target language (in-language) and English (speech-to-English translation, S2ETT).
+S2ETT corresponds to TTest, enabling xlt with monolingual LLMs.
+We use SeamlessM4Tv2-Large and Whisper-v3-Large as ASR backends.
+For languages not supported by the model, we manually select the closest available language for in-language transcription.
+All decoding is greedy, as more complex decoding strategies do not improve xlt performance (Ebing & Glavaš, 2024).
+We additionally evaluate on the original text (Text) of SIB-200 and Belebele.
+For TTest, we translate this text to English using SeamlessM4Tv2-Large.101010We use the gold Q&A in English for TTest.
+Otherwise, MT models would need to be combined, e.g., S2ETT of paragraphs with Whisper, and translation of textual Q&A with SeamlessM4Tv2.
+For classification, we evaluate three text encoders: Roberta-Large (Liu et al., 2019), LLM2Vec (BehnamGhader et al., 2024), and NLLB-LLM2Vec (Schmidt et al., 2024).
+LLM2Vec is a sequence encoder based on Llama 3 8B (AI@Meta, 2024), trained with bidirectional attention on masked next-token prediction and SimCSE (Gao et al., 2021).
+NLLB-LLM2Vec extends LLM2Vec with the encoder of the NLLB translation model, covering 200+ languages for robust multilingual NLU (Schmidt et al., 2024).
+Speech-LLMs.
+We further evaluate two multimodal speech-LLMs, Qwen 2.5 7B-Omni (QwenOmni) and Gemini 2.0 Flash111111https://cloud.google.com/vertex-ai/generative-ai/docs/models/gemini/2-0-flash, in a zero-shot prompting setup.121212Note that the audio encoder of Qwen 2.5 7B-Omni is initialized with Whisper-v3-Large prior to further cross-modal instruction tuning.
+For each task, we provide an English task description and supply the task-specific input (e.g., question, paragraph, or utterance) in the target language, using the appropriate modality (text or speech).
+To ensure consistent audio quality across languages, we normalize all utterances to a root mean square (RMS) level of 0.07.
+For Belebele-Fleurs, where inputs consist of multiple utterances, we normalize the concatenated sequence.
+Appendix §A.1 details the prompt formats and input preprocessing used for each task and modality.
+4.3 Further Details
+Hyperparameters.
+We train all mSE and CS models with AdamW (Loshchilov & Hutter, 2019), weight decay of , and with 10% linear warm-up and followed by linear decay on an effective batch size of 32.
 For SIB-Fleurs, we run a grid search over the learning rates , since suitable hyperparameters have not yet been extensively studied for mSE on such a downstream task.
@@
 We report results for runs on the learning rate that performs best, on average, on the English validation sets.
+For experiments with Qwen 2.5 7B-Omni and Gemini 2-Flash, we perform greedy decoding.
+We report further experimental details of our prompting experiments in §A.1.
 Text vs.
 Utterance Quality.
-We evaluate models trained on the original text (Text) as well as on varying quality of speech.
-For each sentence, Fleurs comprises one or more utterances.
-We compute the character error rate (CER) between the ground truth sentence and its transcriptions.111111We compute CER against the original Flores sentence, as Flores represents the written form of the utterance best suited for processing by LLMs.
-We then fine-tune models on two sets of utterances: those with the lowest CER (best quality) and those with the highest CER (worst quality).
-We thereby isolate the downstream impact of utterance quality, ensure a balanced training set, and compare mSE to CS more fairly, which benefit less from mixing transcriptions of varying quality.
+We evaluate models trained on the original text (Text) as well as on speech data of varying quality.
+Each sentence in Fleurs is associated with one or more utterances.
+To quantify speech quality, we compute the character error rate (CER) between the reference Flores sentence and its transcriptions produced by Whisper-v3-Large and SeamlessM4Tv2-Large.
+Based on these CER scores, we construct two utterance subsets: one comprising the lowest-CER utterances (best quality), and one with the highest-CER utterances (worst quality).
+We fine-tune the mSE and CS models and evaluate all models (incl.
+speech-LLMs) separately on each subset to isolate the effect of utterance quality on downstream performance.
+This setup also balances the training data to enable a fair comparison between mSE, CS, and Text.
 5 Results
-In-Language ZS-XLT Utterance Quality Best Worst Best Worst Best Worst Best Worst Best Worst SIB-Fleurs EN (1) Whisper (85) SeamlessM4T (90) Unsupported (7) Not-EN (101) mSE mHuBERT 41.1 39.7 27.3 19.1 27.2 19.1 26.5 16.3 26.9 18.6 MMS-1b 44.6 46.9 19.9 13.6 19.8 13.5 20.8 13.3 19.8 13.4 MMS-1b-Fleurs 64.8 64.2 25.2 20.8 24.7 20.3 27.1 22.1 25.0 20.6 MMS-1B-All 55.2 53.1 21.7 18.1 21.4 17.9 24.9 21.0 21.7 18.2 Whisper-v3-L 78.3 76.3 46.0 42.0 45.5 41.3 39.4 38.4 44.7 40.8 SeamlessM4Tv2-L 87.4 88.5 82.3 79.3 82.2 79.1 55.7 52.9 79.0 75.8 CS Roberta-L-Wh-EN 90.4 91.7 75.8 72.7 73.6 70.8 52.0 48.7 71.5 68.4 Roberta-L-S4T-EN 91.7 90.8 86.2 84.5 85.9 84.2 56.3 55.2 82.3 80.6 LLM2Vec-L-Wh-EN 91.1 92.3 76.6 75.3 74.5 73.4 53.0 53.2 72.4 71.1 LLM2Vec-L-S4T-EN 93.0 91.3 87.1 85.2 86.9 84.9 58.1 56.6 83.4 81.4 NLLB-LLM2Vec-L-Wh 92.5 91.0 77.5 73.9 75.4 72.2 54.4 53.2 73.7 70.4 NLLB-LLM2Vec-L-S4T 94.0 92.7 85.1 84.9 85.3 84.9 62.3 60.7 82.3 81.9 Text Roberta-L-S4T-EN 92.3 87.4 89.1 55.3 85.2 LLM2Vec-S4T-EN 91.0 86.4 87.9 55.6 84.1 NLLB-LLM2Vec 92.3 88.1 87.9 80.7 87.2 Belebele-Fleurs EN (1) Whisper (79) SeamlessM4T (84) Unsupported (4) Not-EN (91) CS Roberta-L-Wh-EN 82.6 81.3 57.5 56.2 56.4 55.2 46.2 44.7 55.5 54.3 Roberta-L-S4T-EN 82.0 79.4 67.5 66.4 67.0 65.8 47.2 46.2 65.2 64.1 LLM2Vec-Wh-EN 94.9 94.4 74.6 73.5 73.6 72.6 63.5 62.9 72.7 71.8 LLM2Vec-S4T-EN 95.5 94.4 84.0 83.1 83.5 82.5 64.7 63.6 81.8 80.9 NLLB-LLM2Vec-Wh 94.8 93.5 58.0 56.2 56.8 55.0 40.9 41.0 55.5 53.8 NLLB-LLM2Vec-S4T 94.9 93.5 61.3 60.2 60.8 59.6 43.9 41.1 59.2 57.8 Text Roberta-L-S4T-EN 83.3 72.1 72.6 48.3 70.5 LLM2Vec-S4T-EN 95.3 87.8 88.3 66.0 86.3 NLLB-LLM2Vec 95.1 65.6 64.6 55.2 63.8
 Table 3 summarizes the main results for both tasks by approach across our xlt setups. We dissect the results along several axes of analysis.
-English. The first column presents the in-language English performance by task, modality (Text vs. mSE and CS), and utterance quality by model.
-Text & CS.
+English. The first column presents the in-language English performance by task, modality (Text vs. mSE, CS, and speech-LLM), and utterance quality by model.
+Text, CS, and Speech-LLMs.
 English ASR performance is strong for both transcription models (cf.
 Figure 1), and CS effectively leverages the NLU capabilities of LLMs, thanks to their extensive pre-training on both tasks.
-CS are even on par with models trained on gold text, underscoring the quality of ASR on English.
+CS are even on par with and sometimes outperform models trained on gold text, underscoring the quality of ASR on English.
+The slight outperformance of CS over Text in SIB-Fleurs results from model selection on comparatively small validation splits (cf.
+Table 2).
 As a result, CS outperforms all mSE models on SIB-Fleurs.
@@
 In sum, the transcription model has no clear impact on English performance.
+The speech-LLMs perform on par or better on Belebele-Fleurs while slightly trailing the best fine-tuned models on both CS and Text on SIB-Fleurs.
+Gemini 2.0 Flash consistently performs better than Qwen 2.5B 7B-Omni.
+Notably, their cross-modal gap (Text vs.
+speech-LLM) is small (approx.
+-1.5%).
+English Non-English Utterance Quality (Best; Worst) B W B W B W B W B W Language Groups (Size) Setup EN Whisper S4T Unsup.
+Non-EN SIB-Fleurs (1) (85) (90) (7) (101) mSE mHuBERT X 41.1 39.7 27.3 19.1 27.2 19.1 26.5 16.3 26.9 18.6 MMS-1b X 44.6 46.9 19.9 13.6 19.8 13.5 20.8 13.3 19.8 13.4 MMS-1b-Fleurs X 64.8 64.2 25.2 20.8 24.7 20.3 27.1 22.1 25.0 20.6 MMS-1b-All X 55.2 53.1 21.7 18.1 21.4 17.9 24.9 21.0 21.7 18.2 Whisper-v3-L X 78.3 76.3 46.0 42.0 45.5 41.3 39.4 38.4 44.7 40.8 SeamlessM4Tv2-L X 87.4 88.5 82.3 79.3 82.2 79.1 55.7 52.9 79.0 75.8 CS Roberta-Wh-EN T 90.4 91.7 75.8 72.7 73.6 70.8 52.0 48.7 71.5 68.4 Roberta-S4T-EN T 91.7 90.8 86.2 84.5 85.9 84.2 56.3 55.2 82.3 80.6 LLM2Vec-Wh-EN T 91.1 92.3 76.6 75.3 74.5 73.4 53.0 53.2 72.4 71.1 LLM2Vec-S4T-EN T 93.0 91.3 87.1 85.2 86.9 84.9 58.1 56.6 83.4 81.4 NLLB-LLM2Vec-Wh X 92.5 91.0 77.5 73.9 75.4 72.2 54.4 53.2 73.7 70.4 NLLB-LLM2Vec-S4T P 94.0 92.7 85.1 84.9 85.3 84.9 62.3 60.7 82.3 81.9 Speech- Qwen 2.5 7B-Omni P 87.0 87.6 59.5 58.4 58.3 57.3 45.6 45.6 56.9 55.9 LLM Gemini 2.0 Flash P 87.6 88.7 82.1 81.4 80.8 80.0 59.4 59.3 78.6 77.8 Text Roberta-S4T-EN T 92.3 87.4 89.1 55.3 85.2 LLM2Vec-S4T-EN T 91.0 86.4 87.9 55.6 84.1 NLLB-LLM2Vec X 92.3 88.1 87.9 80.7 87.2 Qwen 2.5 7B-Omni P 88.7 76.4 75.1 53.9 73.0 Gemini 2.0 Flash P 90.4 87.0 86.8 77.2 86.1 Belebele-Fleurs (1) (79) (84) (4) (91) CS Roberta-Wh-EN T 82.6 81.3 57.5 56.2 56.4 55.2 46.2 44.7 55.5 54.3 Roberta-S4T-EN T 82.0 79.4 67.5 66.4 67.0 65.8 47.2 46.2 65.2 64.1 LLM2Vec-Wh-EN T 94.9 94.4 74.6 73.5 73.6 72.6 63.5 62.9 72.7 71.8 LLM2Vec-S4T-EN T 95.5 94.4 84.0 83.1 83.5 82.5 64.7 63.6 81.8 80.9 NLLB-LLM2Vec-Wh X 94.8 93.5 58.0 56.2 56.8 55.0 40.9 41.0 55.5 53.8 NLLB-LLM2Vec-S4T X 94.9 93.5 61.3 60.2 60.8 59.6 43.9 41.1 59.2 57.8 Speech- Qwen 2.5 7B-Omni P 91.6 91.0 51.4 50.8 50.4 49.7 34.4 34.4 49.0 48.4 LLM Gemini 2.0 Flash P 94.1 93.4 84.4 83.5 82.9 82.0 59.7 59.4 81.1 80.2 Text Roberta-S4T-EN T 83.3 72.1 72.6 48.3 70.5 LLM2Vec-S4T-EN T 95.3 87.8 88.3 66.0 86.3 NLLB-LLM2Vec X 95.1 65.6 64.6 55.2 63.8 Qwen 2.5 7B-Omni P 94.1 69.1 67.4 36.6 64.9 Gemini 2.0 Flash P 95.9 90.8 89.7 82.0 89.1
 mSE.
@@
 The large-scale pre-training of Whisper-v3 on both multilingual ASR and S2ETT enhances its SLU performance.
-SeamlessM4Tv2-Large is the best performing mSE on the English test set of SIB-Fleurs (87.4%), with only a slight performance gap compared to CS (ca.
+SeamlessM4Tv2-Large is the best performing mSE on the English test set of SIB-Fleurs (87.4%), with only a slight performance gap compared to CS (approx.
 -3 pp).
@@
 Utterance Quality.
-The quality of English utterances does not affect performance for either mSE or CS across both tasks.
-This can be attributed to the exhaustive training of all models on English ASR due to the scale of English speech training data.
+The quality of English utterances does not affect performance of the models across both tasks.
+This can be attributed to the large-scale training on English ASR data of all models.
 This is why CS are also on par with Text.
@@
 For each task, we group the languages into (i) languages supported by Whisper-v3, (ii) languages supported by SeamlessM4Tv2, and (iii) languages unsupported by either model.
-Whisper-v3 and SeamlessM4Tv2 overlap in their support for 81 Fleurs languages.121212For unsupported languages, we hand-select the closest supported language to transcribe into.
-Text & CS.
+Whisper-v3 and SeamlessM4Tv2 overlap in their support for 81 Fleurs languages.131313For unsupported languages, we hand-select the closest supported language to transcribe into.
+Text, CS, and Speech-LLMs.
 Across both tasks and all language groups, models trained on transcriptions of SeamlessM4Tv2-Large consistently outperform models fine-tuned on transcriptions of Whisper-v3-Large.
 The performance gap grows as Whisper’s support decreases or becomes unavailable for the target languages (cf.
-‘SeamlessM4Tv2’ and ‘unsupported’ language groups in Table 3).
-Nevertheless, while CS are mostly competitive on SIB-Fleurs, they trail models evaluated on ground-truth paragraphs more significantly on Belebele-Fleurs (ca.
+‘S4T’ and ‘Unsup.’ language groups in Table 3).
+Nevertheless, while CS are mostly competitive on SIB-Fleurs, they trail models evaluated on ground-truth paragraphs more significantly on Belebele-Fleurs (approx.
 -5 pp).
 We presume that transcription and S2ETT errors propagate more severely in multiple-choice QA.
-For SIB-Fleurs, all CS with SeamlessM4Tv2-Large deteriorate only slightly in xlt performance to all 101 target language, on average, relative to English (ca.
+Moreover, the best speech-LLM, Gemini 2.0 Flash, performs highly competitively across tasks, boasting on par or better performance than other models.
+For SIB-Fleurs, all CS with SeamlessM4Tv2-Large deteriorate only slightly in xlt performance to all 101 target language, on average, relative to English (approx.
 -10 pp).
-Most of this xlt gap comes from the 7 target languages that SeamlessM4Tv2 does not support (ca.
+Most of this xlt gap comes from the 7 target languages that SeamlessM4Tv2 does not support (approx.
 -35 pp on average).
-In contrast, pairing LMs with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (ca.
--13 pp) and again a larger deficit on languages neither model supports (ca.
+In contrast, pairing LMs with Whisper-v3-Large causes more pronounced drops on languages that Whisper supports (approx.
+-13 pp) and again a larger deficit on languages neither model supports (approx.
 -32 pp).
-CS are highly competitive to Text (ca.
+CS are highly competitive to Text (approx.
 -0.7 pp).
-Only NLLB-LLM2Vec evaluated on gold text tremendously outperforms all other models on unsupported languages, since they are supported by NLLB Team et al.
-(2022).
+Only NLLB-LLM2Vec evaluated on gold text tremendously outperforms all other models on unsupported languages, since they are supported by NLLB (Team et al., 2022).
 For Belebele-Fleurs, S2ETT paired with LLM2Vec (i.e., TTest) outperforms zs-xlt cascading on in-language transcriptions and NLLB-LLM2Vec, except for unsupported languages.
@@
 Figure 1).
+The performance of speech-LLMs varies substantially across models and language groups.
+Gemini 2.0 Flash strongly outperforms Qwen 2.5 7B-Omni on both tasks in non-English settings.
+Qwen exhibits a marked performance degradation when transitioning from English to non-English inputs, dropping by approx.
+30 pp in Text, and even more steeply in the speech modality, where the cross-modal gap widens from approx.
+3% (English) to 18% (non-English).
+In contrast, Gemini maintains strong performance on non-English Text inputs and shows only a moderate cross-modal drop of approx.
+9%.
+This however exceeds the gap for the best fine-tuned CS models vs.
+their Text counterparts (approx.
+3%).
+Qwen relies on Whisper-v3-Large as its audio encoder, making it sensitive to Whisper’s transcription limitations.
+Moreover, its instruction tuning possibly is focused on a small subset of high-resource languages, potentially limiting its generalization to languages not seen during alignment or fine-tuning (QwenOmni).
+This is consistent with its sharp drop in performance on unsupported languages.
+By contrast, Gemini’s closed-source nature precludes detailed analysis, but its robustness across modalities and languages suggests a more extensive and diverse instruction tuning setup.
+It may also benefit from large-scale multilingual pretraining, enhanced alignment objectives, or broader training coverage.
 mSE.
 The zs-xlt performance of mSE on SIB-Fleurs mirrors the trends we observed for English performance.
-When trained solely with the wav2vec-BERT 2.0 objective, MMS-1B (19%) only slightly surpasses random performance (14%) in zs-xlt for the highest-quality utterances, regardless of the target language group.
-In contrast to English, post-hoc modular ASR fine-tuning of MMS-1B with language adapters and language-specific decoding heads, whether on Fleurs alone (MMS-1B-Fleurs) or on Fleurs with additional data (MMS-1B-All), yields much smaller gains in zs-xlt (ca.
+When trained solely with the wav2vec 2.0 objective, MMS-1B (19%) only slightly surpasses random performance (14%) in zs-xlt for the highest-quality utterances, regardless of the target language group.
+In contrast to English, post-hoc modular ASR fine-tuning of MMS-1B with language adapters and language-specific decoding heads, whether on Fleurs alone (MMS-1B-Fleurs) or on Fleurs with additional data (MMS-1B-All), yields much smaller gains in zs-xlt (approx.
 +1-3 pp).
 Furthermore, despite employing fully parameter-shared multilingual ASR and S2ETT pre-training, Whisper-v3-Large fails to transfer strong English in-language performance to other languages effectively, with the xlt gap ranging from -32.3 pp for supported languages to -38.9 pp for unseen languages.
-In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp) though also deflates in performance on unsupported languages (55.7%).
+In contrast, only SeamlessM4Tv2-Large achieves performance comparable to English on supported languages (-5.2 pp), though also deflates in performance on unsupported languages (55.7%).
 Overall, these findings nevertheless indicate that multilingual cross-modal translation and multilingual text-to-speech distillation align and semantically enrich multilingual speech representations to enhance cross-lingual SLU.
 Utterance Quality.
-The utterance quality consistently affects performance on both tasks for all model configurations in zs-xlt.
-Notably, CS seem to be less affected than mSE.
-We attribute this observation to the much more sizable pre-training on text of diverse quality of LLMs compared to speech models.
+The utterance quality consistently affects performance on both tasks for all model configurations evaluated on non-English languages.
+Notably, CS and speech-LLMs seem to be less affected than mSE.
+We attribute this to the much more sizable pre-training on text of diverse quality of LLMs compared to speech-only models.
 For SIB-Fleurs, the magnitude of the performance gaps between best and worst quality utterances in zs-xlt mimics how well the speech model backbones perform across both tasks in both CS and end-to-end speech classification.
-SeamlessM4Tv2-Large is more robust to noisy utterances than Whisper-v3-Large, whereas MMS variants suffer from the largest drops (ca.
+SeamlessM4Tv2-Large is more robust to noisy utterances than Whisper-v3-Large, whereas MMS variants suffer from the largest drops (approx.
 5 pp).
@@
 To further understand the underlying factors behind our main results (cf.
-Table 3), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages.
-We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.131313We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large.
-We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model is highly comparable across splits.
+Table 3), we benchmark Whisper-v3-Large and SeamlessM4Tv2-Large on ASR and Speech-to-English-text-translation (S2ETT) on all 102 Fleurs languages in Figure 1.
+We first compute CER and sacreBLEU between the Flores sentences and the in-language transcriptions and S2ETT transcriptions of utterances in Fleurs pooled across all splits, respectively.141414We verify that there is no leakage of dev and test splits of Fleurs in SeamlessM4Tv2-Large and Whisper-v3-Large.
+We observe that the averages and the standard deviations CER and sacreBLEU by task and language for either model are highly comparable across splits.
 We trust that the models have not been trained on the dev and test set of Fleurs.
-We use Huggingface’ evaluate to compute for CER and sacreBELU.
-We then ‘macro-average’ the metrics over all languages.
-Figure 1 charts the results of this analysis.
+We ‘macro-average’ the metrics over all languages.
 ASR.
@@
 SeamlessM4Tv2-Large outperforms Whisper-v3-large across the board in speech-to-English-text-translation and is favored for all but 2 languages.
-For both groups of languages supported by, Whisper-v3-Large and SeamlessM4Tv2-Large, respectively, SeamlessM4Tv2-Large achieves on average about 8 higher sacreBLEU than Whisper-v3-Large.
+For both groups of languages supported by Whisper-v3-Large and SeamlessM4Tv2-Large, respectively, SeamlessM4Tv2-Large achieves on average about 8 higher sacreBLEU than Whisper-v3-Large.
 The gap from SeamlessM4T-Large to Whisper-v3-Large reduces to about 5 sacreBLEU for unsupported languages.
@@
 The more robust ASR and much stronger translation performance of SeamlessM4Tv2-Large results stem from its pre-training.
-SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB translation model Team et al.
-(2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on wav2vec-BERT 2.0 objective.
+SeamlessM4Tv2 first initializes a text encoder and decoder with weights from a pre-trained NLLB translation model (Team et al., 2022) and a speech encoder pre-trained on 4.5M hours of self-supervised training on w2v-BERT 2.0 objective.
 The model is then trained on translation objectives from text and speech to text between any two languages in both translation directions.
@@
 The sequence length of encoded speech typically far exceeds the length of embedded tokenized text for the same input.
-The mSE of SeamlessM4Tv2-Large thus appends a temporal convolution, a ‘length adaptor’, as its final layer to reduce the resolution of speech frames by a factor of 8 and to better align the modalities Seamless Communication et al.
-(2023).141414For architectural details on the length adaptor, we refer to the original paper Communication et al.
-(2023).
+The mSE of SeamlessM4Tv2-Large thus appends a temporal convolution, a ‘length adaptor’, as its final layer to reduce the resolution of speech frames by a factor of 8 and to better align the modalities (Seamless Communication et al., 2023).151515For architectural details, we refer to the original paper (Communication et al., 2023).
 This begs the question whether a length adaptor is essential for pooling semantics from speech output tokens.
@@
 We conclude that the pre-training regime is more vital for multilingual SLU capabilities of mSE than nuanced architectural design choices.
-ZS-XLT EN (1) S4T (89) N/A (12) AVG (101) S4Tv2-L Size B W B W B W B W Incl.
-LA 635M 87.4 88.5 82.2 79.1 55.7 52.9 79.0 75.8 Excl.
-LA 588M 88.5 86.4 80.9 77.5 55.6 52.0 77.8 74.4 47M +1.1 -2.1 -1.3 -2.6 -0.1 -0.9 -1.2 -1.4
+We leave a more detailed investigation into the utility of length adaptors in SLU to future work.
+ZS-XLT | |||||||||
+EN (1) | S4T (89) | N/A (12) | AVG (101) | ||||||
+S4Tv2-L | Size | B | W | B | W | B | W | B | W |
+Incl. LA | 635M | 87.4 | 88.5 | 82.2 | 79.1 | 55.7 | 52.9 | 79.0 | 75.8 |
+Excl. LA | 588M | 88.5 | 86.4 | 80.9 | 77.5 | 55.6 | 52.0 | 77.8 | 74.4 |
+47M | +1.1 | -2.1 | -1.3 | -2.6 | -0.1 | -0.9 | -1.2 | -1.4 |
 6 Conclusion
@@
 Our findings indicate that, while cascaded systems remain the most robust option, multilingual speech encoders can achieve competitive performance when adequately pre-trained.
+Moreover, speech-LLMs can yield state-of-the-art performance on par with the best targeted SLU pipelines when appropriately aligned for multilingual spoken language instruction following.
+Moreover, speech-LLMs can achieve state-of-the-art performance on par or better than the best targeted SLU pipelines when properly aligned for multilingual spoken language instruction following.
 Furthermore, we observe a strong correlation between strong multilingual SLU and both the robustness of multilingual ASR and the effectiveness of cross-modal speech translation to English text.
@@
 We hope that our findings inspire future work towards developing more efficient multilingual speech encoders that are jointly pre-trained for both multilingual ASR and SLU to close the performance gap between end-to-end speech classification and cascaded approaches.
-Limitations
-Our work has two primary limitations.
-First, the selection of models in CS can be further expanded.
-In particular, additional evaluations on CS that combine transcription models with prompting 70B LLMs would provide valuable insights given the growing interest in instruction fine-tuned LLMs.
-Moreover, assessing speech-LLMs on both tasks would further illuminate multimodal alignment in massively multilingual scenarios.
-Notably, LLM2Vec on Llama 3.1 8B, which underwent extensive task fine-tuning on 63K instances, represents an exceptionally strong language model backbone for CS.
-Compared to concurrent work, our models would have even outperformed prompting 70B LLMs on speech transcriptions in Belebele Costa-jussà et al.
-(2024).
-Additionally, current speech LLMs rely on Whisper, which is inherently limited to 30-second speech segments.
-This limitation hinders their application in Belebele-Fleurs, where the average speech length is 47 seconds (cf.
-Table 2).
-Our compute infrastructure was unable to support prompting such large LLMs on long contexts, exceeding half a minute of speech combined with corresponding questions and answers.
-We therefore leave these evaluations for future work.
-Furthermore, although our findings indicate a clear link between robust multilingual SLU and strong multilingual ASR, future research must more carefully isolate the mechanisms underlying this interaction.
-Such investigations should clarify the conditions (e.g., training objectives, data and language mixtures) under which SLU most effectively contributes to more robust multilingual ASR.
-Nonetheless, our analysis strongly suggests that multilingual SLU and multilingual ASR are mutually beneficial.
-Acknowledgments
-We use AI assistance, specifically chatGPT o3-mini, to refine both the writing and the tables in our manuscript.
 Appendix A Appendix
@@
 Evaluation, per checkpoint, across all languages supported by the corresponding benchmark, requires about 1 hour due to the comprehensive setups (e.g., original text and two types of transcriptions for speech recordings).
-Code. The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.
+Prompting Qwen 2.5 7B-Omni & Gemini 2-Flash.
+For the multimodal speech-LLMs Qwen 2.5 7B-Omni and Gemini 2-Flash, we apply the following utterance pre-processing.
+Each audio segment is normalized to a target root mean square (RMS) level of 0.07 to ensure consistent loudness across samples and reduce variability due to recording conditions.
+Specifically, we first normalize the loudness of individual utterances; for Belebele-Fleurs, we additionally normalize the loudness after concatenation.
+The models are then prompted using the task- and modality-specific templates detailed below.
 A.2 SIB-Fleurs